### v2

# How Can We Effectively Expand the Vocabulary of LLMs with 0.01GB of Target Language Text?

Abstract
Large language models (LLMs) have shown remarkable capabilities in many languages beyond English. Yet, LLMs require more inference steps when generating non-English text due to their reliance on English-centric tokenizers and vocabulary, resulting in higher usage costs to non-English speakers. Vocabulary expansion with target language tokens is a widely used cross-lingual vocabulary adaptation approach to remedy this issue. Despite its effectiveness in inference speedup, previous work on vocabulary expansion has focused on high-resource settings assuming access to a substantial amount of target language data to effectively initialize the embeddings of the new tokens and adapt the LLM to the target language. However, vocabulary expansion in low-resource settings has yet to be explored. In this paper, we investigate vocabulary expansion in low-resource settings by considering embedding initialization methods and continual pre-training strategies. Through extensive experiments across typologically diverse languages, tasks and models, we establish a set of strategies to perform vocabulary expansion for faster inference, maintaining competitive downstream performance to baselines with only 30K sentences ($\sim$0.01GB text data) from the target language. [^1] Our code and models are available on GitHub.
How Can We Effectively Expand the Vocabulary of LLMs
with 0.01GB of Target Language Text?
Atsuki Yamaguchi1, Aline Villavicencio1,2  and Nikolaos Aletras1 1School of Computer Science, University of Sheffield, United Kingdom 2Department of Computer Science, Institute of Data Science and Artificial Intelligence, University of Exeter, United Kingdom {ayamaguchi1,a.villavicencio,n.aletras}@sheffield.ac.uk
Figure 1: We explore various adaptation strategies for cross-lingual vocabulary expansion (Refer to caption) that allow inference speedups while retaining competitive performance to the source models in low-resource settings.
1 Introduction
Large language models (LLMs) have strong capabilities in English and other languages OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Groeneveld et al. (2024). Yet, processing non-English texts with LLMs is challenging. They suffer from tokenization overfragmentation [^2] See Appendix E.1 for a quantitative analysis. and thus require more inference steps due to the reliance on English-centric tokenizers and vocabulary, resulting in higher utility costs for non-English speakers Ahia et al. (2023); Petrov et al. (2023).
Cross-lingual vocabulary adaptation (CVA) via vocabulary expansion has been proposed to adapt LLMs (including their tokenizers) to specific target languages. Vocabulary expansion approaches extend the vocabulary of a source model with tokens from a target language, followed by continual pre-training on target language data. A wide range of language-specific LLMs derived from an English-centric LLM such as Llama2 Touvron et al. (2023) have been made available following this approach, including Chinese Cui et al. (2023), Tamil Balachandran (2023), Portuguese Larcher et al. (2023), and Japanese Fujii et al. (2024) models, inter alia. Vocabulary expansion improves inference speed but often assumes access to a substantial amount of target language data for adaptation. For example, Chinese and Tamil LLaMAs make use of 20 and 12GB of target language text, respectively Cui et al. (2023); Balachandran (2023). Given the guaranteed large number of model updates in high-resource settings, the embeddings of new tokens are often simply randomly initialized, followed by the standard continual pre-training with a causal language modeling (clm) objective Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024). However, it is not clear how effective this approach to vocabulary expansion is in low-resource settings.
In this paper, we seek to (i) answer if this widely used adaptation approach under high-resource settings is as effective in low-resource settings; and (ii) identify the best possible vocabulary expansion strategies for language adaptation in low-resource settings, maintaining similar performance to the source models with faster inference (Figure 1). Our key contributions are as follows:
  • We present the first systematic study of vocabulary expansion-based adaptation of generative LLMs (i.e three English-centric models) in low-resource settings (i.e. assuming only 30K sentences, $\sim$0.01GB text data), on two generation tasks (i.e. machine translation and summarization) across ten typologically diverse languages.
  • Our results show that the popular vocabulary expansion approach in high-resource settings (i.e. random initialization and fine-tuning the full model with a clm objective) is not always optimal in low-resource settings.
  • We find that vocabulary initialization approaches that use heuristics information from the source and target tokinizers are more effective. Furthermore, fine-tuning the top and bottom two layers of the LLM using a multi-token prediction objective Gloeckle et al. (2024) works better than fine-tuning the full model with clm. Finally, using a short input sequence length by splitting longer text into multiple sentences allows for a larger number of model updates, mitigating underfitting.
2 Related Work
Text Overfragmentation.
LLMs tend to overfragment text in underrepresented languages Rust et al. (2021); Muller et al. (2021). Overfragmentation has significant implications for non-English speakers, including higher API costs Ahia et al. (2023); Petrov et al. (2023), slower inference Hofmann et al. (2022); Sun et al. (2023); Petrov et al. (2023), and lower downstream performance Bostrom and Durrett (2020); Rust et al. (2021); Toraman et al. (2023); Fujii et al. (2023).
Cross-lingual Vocabulary Adaptation.
CVA methods have been proposed for adapting base LLMs to specific target languages for improving downstream performance and inference speed Cui et al. (2023); Balachandran (2023); Fujii et al. (2024); Yamaguchi et al. (2024). CVA via vocabulary expansion incorporates new tokens into the source vocabulary Chau et al. (2020); Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Fujii et al. (2024); Liu et al. (2024). CVA with vocabulary replacement replaces the entire source vocabulary with a new one from the target Minixhofer et al. (2022); Dobler and de Melo (2023); Ostendorff and Rehm (2023); Downey et al. (2023). More recent methods extend only the language modeling head Hong et al. (2024). CVA is followed by continual pre-training on target language data, often called language adaptive pre-training – LAPT Chau et al. (2020).
Vocabulary expansion is widely used for developing language-specific generative LLMs from a source model, e.g. Chinese Cui et al. (2023) and Tamil Balachandran (2023) LLaMAs, and Swallow for Japanese Fujii et al. (2024). This line of work assumes access to a substantial amount of target language data, e.g. 312B characters of Japanese text used for Swallow. This might not be feasible in low-resource settings with limited target language data or compute. To the best of our knowledge, our work is the first to investigate vocabulary expansion for efficient LLM inference in extremely low-resource settings by assuming access to a small amount of target language data.
3 Problem Statement
Let $\mathcal{M}_{\text{s}}$ be a source LLM with a vocabulary $\mathcal{V}_{\text{s}}$. Given an additional target vocabulary $\mathcal{V}_{\text{new}}$, the aim is to learn a model $\mathcal{M}_{\text{t}}$ with the identical architecture as $\mathcal{M}_{\text{s}}$ that supports a target vocabulary $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$. $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{t}}$ have their corresponding tokenizer, $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
First, a target language-specific auxiliary tokenizer $\mathcal{T}_{\text{aux}}$ is trained on target language data $\mathcal{D}$. $\mathcal{T}_{\text{t}}$ is then derived from $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{aux}}$ along with their corresponding vocabularies $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{aux}}$ by incorporating top $k$ frequent tokens ($\mathcal{V}_{\text{new}}$) that are in $\mathcal{V}_{\text{aux}}\backslash(\mathcal{V}_{\text{s}}\cap\mathcal{V}_{\text{aux}})$ into $\mathcal{V}_{\text{s}}$, resulting in $\mathcal{V}_{\text{t}}$. Subsequently, $\mathcal{M}_{\text{t}}$ is initialized with the weights of $\mathcal{M}_{\text{s}}$. Its embedding and output layer matrices are expanded such that the former is of dimensionality $|\mathcal{V}_{\text{t}}|\times H_{\text{t}}$ and the latter $H_{\text{t}}\times|\mathcal{V}_{\text{t}}|$, where $H_{\text{t}}$ is the hidden dimensionality of $\mathcal{M}_{\text{t}}$. Each new token representation is initialized by applying a target vocabulary initialization method (§4). If the weights of both matrices are not tied, they are initialized separately. Finally, $\mathcal{M}_{\text{t}}$ is adapted to the target language by training it on $\mathcal{D}$ using a clm objective.
4 Target Vocabulary Initialization
After vocabulary expansion, the new embeddings should be initialized. Our aim is to investigate the effectiveness and robustness of different initialization approaches in low-resource settings. For that purpose, we evaluate: (1) random initialization; (2) initialization based on auxiliary models; and (3) heuristic-based initialization.
4.1 Random Initialization
We randomly initialize new tokens by sampling from $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma$ are the mean and standard deviation of token embeddings of $\mathcal{M}_{\text{s}}$ (Random). This is the most simple and common approach when adapting English-centric LLMs to a target language via vocabulary expansion in high-resource settings Cui et al. (2023); Balachandran (2023); Larcher et al. (2023); Choi et al. (2024).
4.2 Initialization Based on Auxiliary Models
FOCUS (Dobler and de Melo, 2023) is a state-of-the-art CVA method that relies on auxiliary embeddings, i.e. fastText Bojanowski et al. (2017), for initialization. [^3] Due to resource constraints, we only report results using FOCUS as a representative initialization method that relies on auxiliary embeddings since it outperforms other similar methods such as WECHSEL Minixhofer et al. (2022) for vocabulary replacement. The main assumption is that semantic transfer of embeddings should result in better initialization over Random. We apply FOCUS by tokenizing $\mathcal{D}$ using $\mathcal{T}_{\text{t}}$ and train a fastText model for each language. [^4] Preliminary experiments with FOCUS using off-the-shelf pre-trained word-level fastText models yielded lower performance (see Appendix E.2).
4.3 Heuristic-based Initialization
Sophisticated methods such as FOCUS require auxiliary embeddings trained in the target language, which might not be available or hard to train in low-resource settings. Motivated by this, we evaluate a set of heuristic-based initialization methods that do not rely on any external data or model and can be applied to any language.
Mean.
A straightforward approach is to initialize each new token in $\mathcal{V}_{\text{new}}$ with the average of its corresponding source tokens obtained by using $\mathcal{T}_{\text{s}}$ Yao et al. (2021). Koto et al. (2021) and Gee et al. (2022) have also followed a similar approach for vocabulary replacement in domain adaptation.
Merge.
Mean uses solely $\mathcal{T}_{\text{s}}$, which might produce subtokens from $\mathcal{V}_{\text{s}}$ that are not semantically related to the new target token. To overcome this issue, we propose using merge rules from $\mathcal{T}_{\text{t}}$ to effectively initialize $\mathcal{V}_{\text{new}}$. Merge rules describe how $\mathcal{T}_{\text{t}}$ can combine two subtokens into one, e.g. (‘a’, ‘t’) $\rightarrow$ ‘at’. Therefore, each new token in $\mathcal{V}_{\text{new}}$ can be decomposed into several existing subtokens from $\mathcal{V}_{\text{s}}$. We hypothesize that such subtokens can be more semantically related to the new token than those in Mean as they use information from $\mathcal{T}_{\text{t}}$.
We first identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$. Subsequently, we generate a mapping from each new target token to source subtokens using the identified merge rules. Finally, for each new token, we compute the hierarchical mean of the associated source subtoken embeddings using the mapping information. Algorithm 1 in Appendix §A describes our method.
Align.
The tokenization of a given word can change dynamically depending on the context, e.g. ‘the’ can be tokenized as (‘th’, ’e’) or (‘t’, ‘he’). Likewise, a new token in $\mathcal{V}_{\text{t}}$ is not always tokenized in the same way by $\mathcal{T}_{\text{s}}$ when operating on a sequence rather than a single token, which is the case for Mean. To obtain a more fine-grained semantic representation of a new token, we propose token alignment initialization that leverages mapping information between tokens tokenized with $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$ and the mapping frequency (i.e. more frequent mappings are more important for the final representation). This process allows us to consider different tokenization variants for the initialization of a given token.
For each sentence $x\in\mathcal{D}$, we first tokenize it using $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$. We then compare the two tokenized versions and generate a mapping list of tuples for each new token in $x$, e.g. a token $t_{5}$ in $\mathcal{V}_{\text{t}}$ is tokenized as [($t_{1}$, $t_{2}$), ($t_{3}$, $t_{4}$)] in $\mathcal{V}_{\text{s}}$. Subsequently, we concatenate all mapping lists of tuples in $\mathcal{D}$ for each new token $t\in\mathcal{V}_{\text{new}}$. For each $t$, we then construct a unique set of its constituent tuples, i.e. $\bm{T}_{t}=(\bm{t}_{1},\bm{t}_{2},\bm{t}_{3},\dots)$, and the corresponding frequency list $\bm{F}_{t}=(f_{\bm{t}_{1}},f_{\bm{t}_{2}},f_{\bm{t}_{3}},\dots)$, given the concatenated list in the previous step. Finally, we use $\bm{T}_{t}$ and $\bm{F}_{t}$ to aggregate the mapping information for token $t$ and generate its representation by computing $\sum_{\bm{t}\in\bm{T}_{t}}{\left[f_{\bm{t}}\frac{1}{|\bm{t}|}\sum_{t^{\prime}\in\bm{t}}{\bm{e}_{t^{\prime}}}\right]}$, where $\bm{e}_{t^{\prime}}$ is the vector representation (embedding or LM head) of the subtoken $t^{\prime}\in\mathcal{V}_{s}$.
5 Training Strategy
Continual pre-training on $\mathcal{D}$ is an integral part of vocabulary expansion to improve the alignment of the newly initialized embeddings. To this end, we explore: (i) the training procedure, (ii) the objective function, and (iii) the input sequence length.
Procedure.
We use by default low-rank adaptation (LoRAHu et al. (2022) applied to all linear layers, following previous work on vocabulary expansion in high-resource settings Cui et al. (2023); Balachandran (2023); Abbasi et al. (2023); Choi et al. (2024).
We also consider the two-stage tuning process (2-stage) Cui et al. (2023), where only the embeddings and LM head are first updated, followed by tuning the LoRA modules. This process can help minimize the risk of overfitting to the initial embedding state, which might be suboptimal Downey et al. (2023). Alternatively, we train only the top and bottom two layers (2x2 LS) following Remy et al. (2024). This approach aims to calibrate only the parts closely related to the encoding and decoding of the target language, minimizing changes to the base model. Note that none of these approaches have ever been compared to the standard LoRA or against each other. We tune the embeddings and LM head in each case.
Objective.
We first evaluate clm, which has been used by previous work in high-resource settings (Cui et al., 2023; Balachandran, 2023; Larcher et al., 2023, inter alia), as a default training objective. We also consider a multi-token prediction (mtp) objective Gloeckle et al. (2024), where the model must predict multiple consecutive tokens at each timestep rather than a single token. mtp has been proposed for pre-training and exhibited performance gains over clm. However, it has yet to be explored for continual pre-training for cross-lingual transfer. We experiment with one additional LM head (i.e. predicting two consecutive tokens at a time) and initialize the additional weights with those from the original LM head. [^5] We find in our preliminary analysis that random initialization does not work well (Appendix E.3).
Sequence Length.
Finally, we shorten the default input sequence length from 2,048 to 512, thereby increasing the number of training batches. We hypothesize that this is critical in low-resource settings, as a small number of model updates could be prone to underfitting.
6 Experimental Setup
6.1 Source Models
We use Llama2 (7B) Touvron et al. (2023), an English-centric model, with its $\mathcal{T}_{\text{s}}$ based on byte-fallback Byte Pair Encoding (BPE) Sennrich et al. (2016) and $|\mathcal{V}_{\text{s}}|$ set to 32K in the experiments. We also use Llama3 (8B) Dubey et al. (2024) and Gemma2 (9B) Riviere et al. (2024) for consistency in our analysis. These models have a far larger vocabulary size than Llama2, i.e. 128K and 256K, respectively.
6.2 Target Languages and Data
We experiment with a typologically diverse set of ten languages with various scripts. This includes German (Indo-European) and Swahili (Niger–Congo) for the Latin script, and Arabic (Afroasiatic), Burmese (Sino-Tibetan), Greek (Indo-European), Hindi (Indo-European), Japanese (Japonic), Sinhala (Indo-European), Telugu (Dravidian), and Thai (Kra–Dai) for non-Latin scripts. We select these languages because of the availability of downstream task datasets with the same task formulation across languages.
To simulate a realistic low-resource adaptation scenario, we follow Yong et al. (2023) in using 30K sentences per language. These sentences are randomly sampled from their language-specific subcorpus of CC-100 Conneau et al. (2020).
Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
Afroasiatic Sino-Tibetan Indo-European Indo-European Indo-European Japonic Indo-European Niger–Congo Dravidian Kra–Dai
Model mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum
Source .08.03 360.0 .03.00 240.1 .25.01 250.1 .08.03 250.2 .12.03 390.2 .03.03 220.1 .05.00 260.1 .01.00 270.4 .06.00 210.2 .06.02 220.2
LAPT .18.00 360.2 .11.00 240.3 .20.00 210.1 .21.00 250.3 .21.00 380.2 .08.00 230.2 .07.00 274.1 .12.00 300.1 .09.00 290.1 .12.00 230.2
+Speedup Random .07.00 320.3 .05.00 170.3 .28.00 230.3 .13.00 200.1 .09.00 350.2 .18.00 220.2 .07.00 290.1 .13.00 290.1 .06.00 260.2 .07.00 190.2
FOCUS .07.00 320.2 .03.00 160.3 .28.00 220.4 .17.00 180.1 .10.00 340.2 .19.00 220.2 .05.00 290.3 .18.00 290.1 .05.00 260.2 .06.00 170.1
Mean .06.00 330.1 .04.00 200.4 .24.01 230.2 .14.00 190.1 .12.00 340.2 .20.00 230.1 .06.00 300.3 .12.00 290.1 .06.00 270.1 .09.00 220.1
Merge .07.00 330.1 .04.00 90.5 .25.00 230.2 .15.00 180.3 .10.00 340.2 .18.00 230.3 .09.00 300.2 .13.00 290.1 .05.00 260.3 .07.00 210.2
Align .06.00 330.1 .04.00 150.7 .26.01 210.2 .16.00 170.2 .13.00 350.2 .20.00 230.1 .08.00 300.3 .17.00 290.0 .06.00 270.2 .07.00 220.2
Table 1: Mean performance on mt and sum tasks over five runs in low-resource settings (30K sentences) using Llama2 as source. The baselines without vocabulary expansion are in gray. Darker blue and red shades indicate higher positive and negative relative performance change over Source per language and task, respectively.
6.3 Baselines
We use the following two methods as our baselines: (1) the source model $\mathcal{M}_{\text{s}}$ off-the-shelf without any adaptation (Source); (2) we continue pre-training Source on $\mathcal{D}$ by keeping $\mathcal{V}_{\text{s}}$ (LAPT). These do not offer any inference speedups.
6.4 Tasks
We evaluate all models using two tasks per target language in a zero-shot setting: (1) English-to-target machine translation (mt) using FLORES-200 NLLB Team et al. (2022) and (2) summarization (sum) including German MLSUM Scialom et al. (2020), GreekSUM Evdaimon et al. (2023), and XL-Sum Hasan et al. (2021) for the rest.666sum performance is not directly comparable between different languages as the data does not match across languages. Following Ahia et al. (2023), we use 500 random samples. The prompt templates can be found in Appendix B.
6.5 Evaluation Metrics
Task Performance.
We use chrF Popović (2015) for mt and ROUGE-L Lin (2004) for sum. We report average performance over five different runs throughout the experiments and analysis.
Perplexity.
We report perplexity on the 100K language-specific held-out sentences sampled from the CC-100 subcorpus as an auxiliary metric for evaluating model performance.
Inference Efficiency.
We measure inference efficiency as the number of tokens generated per second Hong et al. (2024).
6.6 Implementation Details
We set $|\mathcal{V}_{\text{aux}}|$ to 50K across languages and the number of new target tokens $k$ to 100 by default. We investigate the effect of varying $k$ in §7.4. For LAPT, we train each model for two epochs with a batch size of 8, a maximum learning rate of 1e-4, and a sequence length of 2,048. We set a LoRA rank to 8 following previous work Cui et al. (2023); Abbasi et al. (2023); Lin et al. (2024).
7 Results and Analysis
7.1 Target Vocabulary Initialization
We analyze the effect of different target vocabulary initialization methods (§4) on task performance and inference efficiency using Llama2 as source.
Model ar my de el hi ja si sw te th
Source 8.3 5.0 35.7 4.8 6.4 20.4 3.5 47.2 2.4 9.4
LAPT 4.2 2.7 10.9 2.9 3.2 5.4 2.3 12.7 1.8 4.3
Random 11.9 11.9 12.0 7.1 8.3 15.0 8.7 13.9 7.1 8.8
FOCUS 11.5 13.8 12.1 6.6 8.9 14.8 9.3 13.7 7.9 9.4
Mean 9.4 11.6 11.7 6.0 6.4 14.7 8.1 13.5 6.9 7.8
Merge 9.8 12.6 11.8 6.1 6.7 15.1 8.7 13.7 7.3 8.0
Align 9.3 11.2 11.7 5.9 6.3 15.1 8.0 13.5 6.7 7.6
Table 2: Perplexity on language-specific held-out dataset using Llama2 as source. Note that results are not comparable between models with gray and others due to their difference in vocabulary. Bold and underlined indicate the best and second-best perplexities among adapted models for each language.
Model Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum mt sum
42.2 20.1 28.6 15.0 42.6 24.8 42.4 13.3 42.2 21.5 42.8 21.9 28.1 14.8 43.0 24.1 27.7 10.0 42.6 17.9
Source 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x
LAPT 1.01x 1.13x 0.98x 1.01x 0.97x 1.05x 1.00x 0.97x 1.00x 1.01x 0.98x 1.03x 1.00x 1.00x 0.98x 1.21x 0.98x 0.99x 0.98x 1.01x
Random 1.17x 1.88x 1.18x 1.48x 0.95x 0.97x 1.32x 1.79x 1.19x 1.57x 1.04x 1.13x 1.56x 2.54x 0.92x 1.27x 1.93x 4.26x 1.06x 1.55x
FOCUS 1.14x 1.79x 1.13x 1.38x 0.89x 0.94x 1.41x 1.66x 1.11x 1.49x 1.04x 1.11x 1.13x 2.40x 0.93x 1.26x 1.79x 4.26x 1.06x 1.44x
Mean 1.08x 1.99x 1.02x 2.15x 0.95x 0.95x 1.32x 1.73x 1.15x 1.70x 1.08x 1.13x 1.29x 2.67x 0.91x 1.25x 1.84x 4.58x 1.10x 1.62x
Merge 1.17x 1.95x 1.08x 1.20x 0.94x 1.00x 1.40x 1.70x 1.07x 1.57x 1.05x 1.12x 1.58x 2.49x 0.92x 1.25x 1.71x 4.38x 1.05x 1.59x
Align 1.10x 1.98x 1.08x 1.90x 0.95x 1.04x 1.37x 1.67x 1.15x 1.77x 1.09x 1.13x 1.47x 2.58x 0.93x 1.26x 1.84x 4.42x 1.08x 1.63x
Table 3: Inference efficiency measured by the number of tokens generated per second (T/S). The first row of Source corresponds to its respective T/S for each language and task. Bold and underlined indicate the best and second-best speedups across tasks and models with the same base model for each language.
Task Performance and Perplexity.
Table 1 shows the performance of all methods, while Table 2 shows the corresponding perplexities on the held-out language-specific dataset. We first find that models adapted with Mean and Align show comparable or better performance than Source in the majority of the cases. Specifically, they outperform or perform the same as Source in 13 and 14 out of 20 cases, respectively. This result aligns with the perplexity results in Table 2, where Align tends to exhibit the lowest perplexity across languages, followed by Mean. Further, some performance gaps in the rest of the cases are within only a few points ($\leq$ 3). In contrast, Merge shows worse performance than Align and Mean, outperforming or performing the same as Source in 11 cases. We speculate that Merge might be less informative than these two methods, as it does not rely on surface information for initialization, resulting in lower performance.
Random performs the same as or better than Source in 12 cases, while FOCUS achieves this in 11 cases—the worst of the adapted models along with Merge. This is also evident in the perplexity results, where FOCUS shows the worst perplexity in six languages (i.e. my, de, hi, si, te, and th). These results suggest that sophisticated initialization approaches do not always work well, and the popular Random is not the best choice in low-resource settings. Thus, we conclude that models adapted with Mean and Align are more likely to perform competitively to Source.
Next, we analyze the results by language and task. We see that the adapted models, especially Mean and Align, generally perform competitively with Source within a few points of degradation across all languages and substantially outperform it in Greek, Japanese and Swahili mt, with up to 17 points of improvement in Japanese. However, they often do not perform competitively in Burmese, German, Greek, and Hindi sum. We hypothesize that sum might require more training data than mt to perform well as generating long text (up to 128 generated tokens and using long context) requires models to have strong generative capabilities in a target language. We later address this challenge in §7.2 by showing that these performance gaps can be drastically narrowed using our alternative training strategies, further contributing to the competitiveness of vocabulary expansion approaches to Source.
Inference Efficiency.
Table 3 shows the inference efficiency of adapted models across tasks and languages. We first see that FOCUS tends to exhibit the worst speedups across languages in sum, followed by Merge. Specifically, its speedups are the worst in 8 out of 10 languages. These results are consistent with downstream performance and may be due to an inability to effectively generate newly added target tokens, which is essential for improving inference efficiency in a target language. For mt, we generally see smaller differences between models since its input is mostly in English, and thus, only the output contributes to inference speedups.
Looking at the results by language and script, we observe that non-Latin script languages, except for Japanese, show substantial speedups of up to 4.58x. Japanese, on the other hand, exhibits moderate inference speedups of up to 1.13x. This is possibly due to the inclusion of Japanese (0.1%) and Chinese [^7] Japanese shares some scripts (i.e. kanji) with Chinese. data (0.13%) in the Llama2 training corpus Touvron et al. (2023). For a similar reason, German, which accounts for the largest size among non-English languages in the corpus (0.17%), shows no speedups across tasks. Although Swahili is not included in the training corpus, it does not seem to benefit well from vocabulary expansion with Llama2, possibly due to its Latin script (91% of the pre-training data).
Recommendation.
Using Mean or Align can provide better downstream performance and inference speedups. Vocabulary expansion contributes well to inference speedups when a target language is not included in pre-training data and is written in non-Latin scripts. [^8] If inference speedups do not matter, we recommend using LAPT in general. Since it does not involve parameter updates from scratch, it can lead to more stable task performance in low-resource settings.
7.2 Training Strategy
We analyze the effectiveness of each training strategy introduced in §5. Due to limited compute, we only use the best-performing Align as the vocabulary initialization method and experiment with six languages with the largest speedups (i.e. ar, my, el, hi, si, and te) in sum. Table 4 lists the performance of the adapted models.
Arabic Burmese Greek Hindi Sinhala Telugu
Model mt sum mt sum mt sum mt sum mt sum mt sum
Source .08.03 360.0 .03.00 240.1 .08.03 250.2 .12.03 390.2 .05.00 260.1 .06.00 210.2
LAPT .18.00 360.2 .11.00 240.2 .21.00 250.3 .21.00 380.2 .07.00 274.0 .09.00 290.1
LoRA clm+2048 .06.00 330.1 .04.00 150.7 .16.00 170.2 .13.00 350.2 .08.00 300.3 .06.00 270.2
mtp+2048 .11.00 320.2 .03.00 140.7 .14.00 200.2 .12.00 340.2 .08.00 290.1 .04.00 250.1
clm+512 .12.00 330.1 .12.00 230.1 .23.00 190.4 .17.00 350.1 .11.00 320.1 .10.00 280.1
mtp+512 .14.00 330.1 .12.00 230.1 .21.00 210.3 .15.00 350.1 .11.00 300.3 .09.00 280.1
2-stage clm+2048 .09.00 330.1 .04.00 170.2 .13.00 210.1 .13.00 360.1 .09.00 300.4 .05.00 280.1
mtp+2048 .13.00 330.1 .02.00 170.3 .19.00 210.1 .11.00 360.1 .08.00 290.2 .03.00 250.2
clm+512 .12.01 330.1 .07.00 130.5 .22.00 220.3 .16.00 350.1 .10.00 310.3 .07.00 280.1
mtp+512 .11.00 330.2 .07.00 170.2 .21.00 220.1 .16.00 350.1 .08.00 310.3 .07.00 280.1
2x2 LS clm+2048 .11.01 330.1 .14.00 250.1 .24.00 220.3 .15.00 360.2 .11.00 320.2 .11.00 290.1
mtp+2048 .15.01 330.1 .11.00 230.2 .23.00 230.2 .12.00 360.0 .12.00 320.1 .12.00 280.2
clm+512 .11.00 330.1 .15.00 260.2 .25.00 230.3 .19.00 360.1 .11.00 330.2 .13.00 300.1
mtp+512 .17.00 330.2 .16.00 260.2 .24.00 230.2 .15.00 360.1 .13.00 330.3 .12.00 290.1
Table 4: Mean performance of Align models (five runs) with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Looking at the results by training procedure, we observe that 2x2 LS drastically improves performance across tasks and languages, achieving up to 10 points improvement from LoRA (Burmese mt and sum with clm+2048). This suggests that the frequently used full LoRA approach in high-resource settings is not the best training approach in low-resource settings. Although 2-stage does not consistently provide gains over LoRA, especially for mt, it does not show substantial performance degradation either. Next, we look at the effectiveness of mtp. Although it does not always improve performance over clm, it boosts Arabic mt performance across models, especially with 2x2 LS. Similar to 2-stage, its performance degradation, if it exists, is generally within a few points ($\leq 3$). Finally, usign a short sequence length (512) works well across models and languages, showing improvements of up to 9 points (LoRA+mtp in Burmese and 2-stage+clm in Greek mt), confirming our hypothesis (§5) that increasing the number of model updates can help avoid underfitting.
Recommendation.
2x2 LS generally leads to performance improvements. The short training sequence length of 512 can also contribute to performance improvements in low-resource settings. Finally, mtp sometimes improves performance with no major performance drop in other cases.
7.3 Experiments with Other Source LLMs
We investigate whether models other than Llama2 adapted based on our recommendations in §7.1 and §7.2 benefit from inference speedups while maintaining competitive performance to their base models. We use Llama3 and Gemma2 as source. We apply the 2x2 LS+mtp+512 training strategy and initialize models with either Mean or Align along with Random as a baseline. Due to resource constraints, we only experiment with Burmese, Sinhala, and Telugu in the remainder of the paper, as they are the worst fragmented languages of our target languages (see Appendix E.1). Table 5 shows the performance of adapted models using Llama3 and Gemma2 as source.
Burmese Sinhala Telugu
Llama3 mt sum mt sum mt sum
Source .09.00 280.0 .19.00 300.0 .24.00 290.0
LAPT .17.00 280.1 .21.00 310.1 .30.00 290.1
+Speedup Random .19.00 230.2 .14.00 310.2 .13.00 290.1
Mean .19.00 230.2 .12.00 320.1 .12.00 290.2
Align .21.00 280.3 .22.00 330.1 .31.00 310.1
Speedup 2.60x 3.52x 2.36x 3.19x 2.15x 3.55x
Gemma2
Source .21.01 340.1 .19.00 340.2 .31.01 300.2
LAPT .28.00 280.3 .29.00 340.1 .43.00 290.1
+Speedup Random .24.00 260.3 .26.00 330.1 .35.00 280.2
Mean .25.00 250.3 .27.00 320.2 .33.00 280.1
Align .25.00 250.2 .28.00 320.1 .32.00 290.1
Speedup 2.24x 1.53x 2.26x 1.53x 1.03x 1.25x
Table 5: Mean performance and inference speedup over five runs with Llama3 and Gemma2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Overall, models adapted with Align generally perform competitively with Gemma2 Source in Sinhala and Telugu sum tasks and often outperform Llama3 Source in both tasks and Gemma2 Source in mt. Further, Align models exhibit substantial inference speedups of up to 3.55x and 2.26x for Llama3 and Gemma2, respectively. These results confirm the versatility of our strategies regardless of the type of base model. Although models adapted with Mean rival Source, they, along with Random, underperform it in Burmese sum and Sinhala and Telugu mt tasks when using Llama3 as source. This can be due to underfitting, i.e. their perplexities are higher than Align (Appendix E.4).
Examining the results by source model and task, we observe that the adapted models with Llama3 and Gemma2 as source show substantially better mt performance than those with Llama2 despite their similar model size. In particular, the Telugu-adapted models initialized with Align with Llama3 and Gemma2 as source obtain 19 and 20 points better performance than those Llama2 counterparts. This trend suggests a successful cross-lingual transfer of the generative capabilities of the base models. However, sum does not seem to follow the trend, i.e. the adapted models with Llama3 and Gemma2 do not outperform those with Llama2. In fact, their semantic-level performances measured by BLEURT Sellam et al. (2020) improve up to 7 points from those with Llama2 (Appendix E.4). Thus, adapted models actually enjoy their base model capabilities to improve performance at the semantic level rather than the surface level measured by ROUGE-L.
Figure 2: Downstream performance and inference speedup in (a) mt and (b) sum across different $|\mathcal{V}_{\text{new}}|$. Red and gray dotted lines denote Source and LAPT.
7.4 Target Vocabulary Size
The extent to which a target language benefits from inference speedups varies across languages, models, and tasks, as observed from Tables 3 and 5. Although larger $|\mathcal{V}_{\text{new}}|$ can lead to faster inference, it comes with the risk of underfitting in low-resource settings. [^9] Given a small and limited $\mathcal{D}$, the larger $|\mathcal{V}_{\text{new}}|$, the less fragmentation with respect to $\mathcal{D}$, the fewer the number of training tokens, and therefore the fewer model updates. To help select an optimal $|\mathcal{V}_{\text{new}}|$ for obtaining substantial speedups while retaining competitive performance to Source, we conduct a cost-benefit analysis with respect to $|\mathcal{V}_{\text{new}}|$. We experiment with $|\mathcal{V}_{\text{new}}|=\{50,100,500,1\text{K},5\text{K}\}$ while previous studies set $|\mathcal{V}_{\text{new}}|$ to over 10K, e.g. 17,953 for Chinese LLaMA. For comprehensiveness, we employ all three source models and follow the same setup as in §7.3. Figure 2 shows the performance changes and corresponding inference speedups with respect to different $|\mathcal{V}_{\text{new}}|$.
We first observe that the larger $|\mathcal{V}_{\text{new}}|$, the worse the task performance across different source models, languages, and tasks in general. Notably, sum appears to be less robust to the changes in $|\mathcal{V}_{\text{new}}|$ than mt across different models and languages. Adapted models underperform Source even at $|\mathcal{V}_{\text{new}}|=500$ in the majority of the cases, while they, especially when initialized with Align, still outperform or rival Source in mt in almost all the cases. This difference can be due to the difficulty of each task as discussed in §7.1, i.e. sum requires more $\mathcal{D}$ than mt to perform well. These suggest that $|\mathcal{V}_{\text{new}}|$ should be set smaller (e.g. $|\mathcal{V}_{\text{new}}|=100$ in our setup) for tasks like sum to be competitive with Source in task performance.
[…]
Recommendation.
Setting $|\mathcal{V}_{\text{new}}|$ between 500 and 1K for mt and 100 for sum can be a suitable starting point to maintain competitive performance to Source while benefiting greatly from inference speedups under low-resource settings.
8 Conclusion
We investigated cross-lingual vocabulary expansion in low-resource settings across vocabulary initialization approaches and training strategies. Our extensive experiments reveal that a widely used approach in high-resource settings is not always optimal in low-resource settings. In contrast, models adapted by our alternative strategies achieve faster inference while rivaling their base models in task performance. We supplement our analysis with specific recommendations for effective vocabulary expansion. In future work, we will compare vocabulary expansion and vocabulary replacement for efficient LM inference in low-resource settings.
Limitations
Tokenizer.
Heuristic-based target vocabulary initialization with Merge assumes the use of a BPE-based tokenizer which is a common choice in recent LLMs, e.g. Gemma2, Llama3, Llama2, Mistral Jiang et al. (2023), inter alia. Experimenting with other tokenizers, such as Unigram Kudo (2018), is beyond the scope of this paper.
Languages.
Our work covers ten linguistically diverse languages, i.e. Arabic, Burmese, German, Greek, Hindi, Japanese, Sinhala, Swahili, Telugu, and Thai, following previous work on CVA that has also tested a similar number of languages. For instance, Downey et al. (2023) tested ten languages, and Minixhofer et al. (2022) tested eight languages. Experimenting with more languages is an interesting avenue for future work but out of the scope of this paper given our limited computing capacity.
Model Size.
Our experiments consider LLMs of 7B, 8B, and 9B parameters. Investigating the performance of vocabulary expansion with larger models would be valuable in future studies. Limited access to computational resources did not allow us to experiment with larger LLMs.
Tasks.
Our experiments consider two generation tasks, i.e. machine translation and summarization, following Hong et al. (2024). Nonetheless, investigating different types of downstream tasks (e.g. reasoning) would be an interesting direction but out of the scope of this paper, given our limited computation and limited language resources available.
Acknowledgments
We would like to thank Miles Williams, Huiyin Xue, and Constantinos Karouzos for the comments. We acknowledge IT Services at the University of Sheffield for the provision of services for high-performance computing.
Appendix
Appendix A Pseudo-code for Merge Rule-based Initialization
Algorithm 1 describes the pseudo-code of vocabulary initialization using merge rules.
Data: $\mathcal{V}_{\text{s}}$, $\mathcal{V}_{\text{new}}$, Merge rules of $\mathcal{T}_{\text{t}}$: $Merges$, Embeddings of $\mathcal{M}_{\text{t}}$: $\bm{E}$
Result: Updated $\bm{E}$
1
/* Identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$/
2 $Merges_{\text{new}}\leftarrow\text{List}[~{}]$;
3 for $m\in Merges$ do
4       $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
5       $t\leftarrow t_{1}+t_{2}$;
6       if $t\in\mathcal{V}_{\text{new}}$ then
7             $AddToMerges_{\text{new}}(m)$;
8
9       end if
10
11 end for
12
/* Generate a mapping from a new token to source subtokens using $Merges_{\text{new}}$/
13 $Maps\leftarrow\text{Dict}[t\in\mathcal{V}_{\text{new}},\text{List}[~{}]]$;
14 while $|Merges_{\text{new}}|\neq 0$ do
15       $UnmapMerges\leftarrow\text{List}[~{}]$;
16       for $m\in Merges_{\text{new}}$ do
17             $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
18             $t\leftarrow t_{1}+t_{2}$;
19             if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
20                   $AddToMaps[t]([t_{1},t_{2}])$;
21
22            else if $t_{1}\in\mathcal{V}_{\text{new}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
23                   if $Maps[t_{1}]\neq\text{List}[~{}]$ then
24                         $AddToMaps[t]([Maps[t_{1}],t_{2}])$
25                  else
26                         $AddToUnmapMerges(m)$
27                   end if
28
29            else if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{new}}$ then
30                   if $Maps[t_{2}]\neq\text{List}[~{}]$ then
31                         $AddToMaps[t]([t_{1},Maps[t_{2}]])$
32                  else
33                         $AddToUnmapMerges(m)$
34                   end if
35
36             else
37                   if $Maps[t_{1}]\neq\text{List}[~{}]\cap Maps[t_{2}]\neq\text{List}[~{}]$ then
38                         $AddToMaps[t]([Maps[t_{1}],Maps[t_{2}]])$
39                  else
40                         $AddToUnmapMerges(m)$
41                   end if
42
43             end if
44
45       end for
46
47      if $|UnmapMerges|=0$ then
48             break;
49
50       end if
51      $Merges_{\text{new}}\leftarrow UnmapMerges$;
52
53 end while
54
/* Obtain vector representation for each new token using $Maps$/
55 for $t\in\mathcal{V}_{\text{new}}$ do
56       $\bm{E}[t]\leftarrow ComputeHierarchicalMean(\bm{E}[Maps[t]])$;
57 end for
58
Algorithm 1 Merge Rule-based Initialization
Appendix B Prompt Templates
We translate the English prompt templates provided by Ahia et al. (2023) for sum with a machine translation API, following Yong et al. (2023). For mt, we formulate a task-specific English prompt, followed by machine translation for each language. Table 8 shows the prompt templates used in our evaluation.
Table 8: Prompt template for each task and language.
Appendix C Implementation Details
C.1 Language Adaptive Pre-training
Table 6 lists the hyperparameters in LAPT.
C.2 Evaluation
sum.
To compute ROUGE-L, we split sentences with an mT5 Xue et al. (2021) tokenizer as preprocessing following Maynez et al. (2023) and subsequently call rouge_scorer [^10] https://github.com/google-research/google-research/tree/master/rouge to compute the metric. To compute BLEURT, we use BLEURT-20, which is available from https://github.com/google-research/bleurt.
Hyperparameters.
Table 7 lists the parameters used during evaluation. To make a fair comparison, we do not conduct any generation parameter tuning and use the same ones across all approaches. For sum, we truncate an article whenever it exceeds the maximum prompt length of 4,096 to avoid the CUDA out-of-memory error.
C.3 Libraries and Hardware
For Llama2 and Gemma2, we train tokenizers using SentencePiece Kudo and Richardson (2018) and convert them into the Hugging Face Tokenizers Moi and Patry (2023) format. For Llama3, we train tokenizers using Hugging Face Tokenizers. We implement our models using PyTorch Paszke et al. (2019), Hugging Face Transformers Wolf et al. (2020) and PEFT Mangrulkar et al. (2022). We preprocess datasets with Hugging Face Datasets Lhoest et al. (2021). For evaluation, we use Hugging Face LightEval [^11] https://github.com/huggingface/lighteval. We use a single NVIDIA A100 (80GB) for all experiments.
C.4 Code and Models
Our code and models are available on https://github.com/gucci-j/lowres-cve.
Appendix D Licenses
This study uses various publicly available models and datasets with different licenses, as detailed below, all of which permit their use for academic research.
D.1 Models
Llama2 is licensed under the Llama 2 Community License Agreement. [^12] https://ai.meta.com/llama/license/ Likewise, Llama3 is licensed under the Llama 3 Community License Agreement. [^13] https://llama.meta.com/llama3/license/ Gemma2 is licensed under the Gemma Terms of Use Agreement. [^14] https://ai.google.dev/gemma/terms
Figure 3: Average number of tokens on the FLORES-200 dev set across languages and models.
D.2 Datasets
FLORES-200 is distributed under CC BY-SA 4.0. XL-Sum is licensed under CC BY-NC-SA 4.0, while MLSUM is distributed under an MIT license. GreekSUM is not made public, but the official script for reproduction is available on https://github.com/iakovosevdaimon/GreekSUM.
Appendix E Results and Analysis
E.1 Overfragmentation
Figure 3 shows the average number of tokens on the FLORES-200 dev set across languages and models. Following the assumption by Ahia et al. (2023) that “language A is more efficiently tokenized than language B if it uses fewer tokens per sentence on average,” [^15] Note that this assumption is only applicable to the FLORES-200 dataset as it is a parallel machine translation corpus. we can clearly see that (1) Burmese, Sinhala, and Telugu are the worst overfragmented languages among the ten target languages across all models, and (2) English is the most efficiently tokenized language.
E.2 Target Vocabulary Initialization
Table 9 compares models adapted with FOCUS using Llama2 as source in Burmese, Sinhala, and Telugu. Subword-fastText is the exact same model as FOCUS in Table 1 and makes use of only $\mathcal{D}$ to train the fastText model. Word-fastText uses a pre-trained language-specific word-level fastText model provided by Grave et al. (2018), i.e. with access to an external model. We observe that models using a word-level pre-trained fastText model consistently underperform those using a subword-level fastText model except for sum in Sinhala. Therefore, we conclude that using a pre-trained word-level fastText model does not help improve performance for models adapted with FOCUS in low-resource settings.
Burmese Sinhala Telugu
Model mt sum mt sum mt sum
Source .03.00 240.1 .05.00 260.1 .06.00 210.2
LAPT .11.00 240.3 .07.00 274.1 .09.00 290.1
FOCUS
Subword-fastText .03.00 160.3 .05.00 290.3 .05.00 260.2
Word-fastText .01.00 50.2 .04.00 290.3 .03.00 240.2
Table 9: Mean performance of models initialized with different variants of FOCUS over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Arabic Greek Hindi
Model mt sum mt sum mt sum
Source .08.03 360.0 .08.03 250.2 .12.03 390.2
LAPT .18.00 360.2 .21.00 250.3 .21.00 380.2
LoRA mtp + 2048 .11.00 320.2 .14.00 200.2 .12.00 340.2
mtp $\ominus$ Copy + 2048 .10.00 320.1 .15.00 200.1 .11.00 350.2
2-stage mtp + 2048 .13.00 330.1 .19.00 210.1 .11.00 360.1
mtp $\ominus$ Copy + 2048 .10.01 330.1 .17.00 220.2 .10.00 360.1
2x2 LS mtp + 2048 .15.01 330.1 .23.00 230.2 .12.00 360.0
mtp $\ominus$ Copy + 2048 .09.00 330.1 .15.00 210.1 .09.00 350.1
Table 10: Mean performance of models initialized with Align over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
Model my si te
Source 3.7 3.5 3.0
LAPT 1.9 2.0 1.8
Random 6.9 7.1 6.1
Mean 6.7 6.8 5.8
Align 6.0 5.5 4.6
(a) Llama3
Model my si te
Source 48.0 59.9 51.9
LAPT 4.6 4.3 5.2
Random 9.0 6.9 6.7
Mean 9.0 6.9 6.6
Align 9.0 6.9 6.6
(b) Gemma2
Table 11: Perplexity on language-specific held-out dataset using Llama3 and Gemma2 as source. Note that results are not comparable between models with gray and others due to their difference in vocabulary.
E.3 Training Strategies
Table 10 briefly compares mtp approaches with and without additional LM head initialization based on the original LM head weights in Arabic, Greek, and Hindi. $\ominus$ indicates that an additional LM head is randomly initialized. We can see that although models without additional LM head initialization based on their original LM head weights show similar performance to those counterparts in sum, they generally do not outperform those in mt. Therefore, we conclude that it is essential to initialize the weights of an additional LM head with those from the original LM head for effective continual pre-training under low-resource settings.
Burmese Sinhala Telugu
Llama2 ROUGE-L BLEURT ROUGE-L BLEURT ROUGE-L BLEURT
Source 240.1 .03.00 260.1 .19.00 210.2 .13.00
LAPT 240.3 .04.00 27.4.1 .25.00 290.1 .22.00
+Speedup Random 240.1 .07.00 320.2 .26.00 290.1 .23.00
Mean 250.1 .06.00 320.1 .25.00 290.1 .23.01
Align 260.2 .08.01 330.3 .26.01 290.1 .24.00
Llama3
Source 28.00 .06.00 30.00 .24.00 29.00 .22.00
LAPT 280.1 .06.00 310.1 .25.00 300.1 .29.00
+Speedup Random 230.2 .05.00 310.2 .27.00 290.1 .24.01
Mean 230.2 .05.00 320.1 .29.01 290.2 .26.00
Align 280.3 .07.00 330.1 .31.00 310.1 .28.00
Gemma2
Source 340.1 .10.00 340.2 .32.00 300.2 .30.00
LAPT 280.3 .10.00 340.1 .32.01 290.1 .26.00
+Speedup Random 260.3 .10.00 330.1 .34.01 280.2 .26.00
Mean 250.3 .09.00 320.2 .33.00 280.1 .27.00
Align 250.2 .10.00 320.1 .33.01 290.1 .28.00
Table 12: Mean sum performance in ROUGE-L and BLEURT over five runs. Darker blue and red indicate higher positive and negative relative performance change over Source.
E.4 Experiments with Other Source LLMs
Perplexity.
Table 11 shows the perplexities measured on the language-specific held-out dataset using Llama3 and Gemma2 as source.
ROUGE-L vs. BLEURT.
Table 12 shows the sum performances measured by ROUGE-L and BLEURT across models when using the 2x2 LS + mtp + 512 strategy and the Align initialization. As discussed in §7.3, while performances at the surface level measured by ROUGE-L do not substantially change across different base models, those at the semantic level measured by BLEURT change, especially between Llama2 and Gemma2.
E.5 Target Vocabulary Size
Figure 4 shows the average target token ratio across models, tasks, and languages with respect to different $|\mathcal{V}_{\text{new}}|$.
Llama2.
[…]
Llama3.
We observe that the ratio of target tokens for outputs plateaus at around $|\mathcal{V}_{\text{new}}|=500$ in both tasks. After that, there are some drops in the target token ratio in mt. Interestingly, the ratio of target tokens for outputs is quite similar to that for inputs in sum. suggesting that adapted models effectively utilize newly added tokens in outputs. To further optimize inference efficiency, we hypothesize that adding target tokens with a larger number of characters might work. We leave this verification for future work.
Gemma2.
[…]
Moreover, the ratio of target tokens is far lower than that of Llama2 and Llama3, indicating that newly added tokens are underutilized in Gemma2. This could possibly be due to byte-fallback BPE used by Gemma2. Considering Llama3 (i) employs byte-level BPE, (ii) shows substantial speedups with vocabulary expansion in Table 5 and Figure 2, (iii) has also a large vocabulary size of 128K, and (iv) has similar fragmentation rates to Gemma2 for the three languages in Figure 3, vocabulary expansion might work better with byte-level BPE rather than byte-fallback BPE. We hypothesize that since byte-level BPE operates on a byte sequence rather than a raw text, it is, therefore, not affected by language scripts. This nature can help to integrate both a new vocabulary and merge rules with the corresponding source ones. We leave a thorough investigation of this hypothesis for future work.
Figure 4: Average target token ratio in input ($\times$) and output ($\bullet$) with respect to $|\mathcal{V}_{\text{new}}|$ across models, languages, and tasks.

### v3

@@
 However, vocabulary expansion in low-resource settings has yet to be explored.
-In this paper, we investigate vocabulary expansion in low-resource settings by considering embedding initialization methods and continual pre-training strategies.
-Through extensive experiments across typologically diverse languages, tasks and models, we establish a set of strategies to perform vocabulary expansion for faster inference, maintaining competitive downstream performance to baselines with only 30K sentences ($\sim$0.01GB text data) from the target language.
-[^1] Our code and models are available on GitHub.
-How Can We Effectively Expand the Vocabulary of LLMs
-with 0.01GB of Target Language Text?
-Atsuki Yamaguchi1, Aline Villavicencio1,2  and Nikolaos Aletras1 1School of Computer Science, University of Sheffield, United Kingdom 2Department of Computer Science, Institute of Data Science and Artificial Intelligence, University of Exeter, United Kingdom {ayamaguchi1,a.villavicencio,n.aletras}@sheffield.ac.uk
-Figure 1: We explore various adaptation strategies for cross-lingual vocabulary expansion (Refer to caption) that allow inference speedups while retaining competitive performance to the source models in low-resource settings.
+In this article, we investigate vocabulary expansion in low-resource settings by considering embedding initialization methods and continual pre-training strategies.
+Through extensive experiments across typologically diverse languages, tasks and models, we establish a set of strategies to perform vocabulary expansion for faster inference, while striving to maintain competitive downstream performance to baselines.
+This is achieved with only 30K sentences ($\sim$0.01GB text data) from the target language.
+[^1] Our code and models are available via GitHub.
 1 Introduction
-Large language models (LLMs) have strong capabilities in English and other languages OpenAI (2023); Touvron et al.
+Figure 1: We address the challenge of effectively expanding vocabulary for LLMs in low-resource settings.
+This is crucial for reducing inference steps when generating non-English text, as LLMs often rely on English-centric tokenizers and vocabulary.
+Our approach explores various adaptation strategies (Refer to caption) to achieve inference speedups while aiming to retain competitive performance.
+Our recommended strategy combines heuristic-based parameter initialization for new tokens with fine-tuning the top and bottom two layers of the model, using a short input sequence length and a multi-token prediction objective Gloeckle et al.
+(2024).
+Large language models (LLMs) have strong capabilities in English and other languages OpenAI (2023); OpenAI et al.
+(2024); Touvron et al.
 (2023); Jiang et al.
 (2023); Groeneveld et al.
+(2024); DeepSeek-AI et al.
 (2025).
 Yet, processing non-English texts with LLMs is challenging.
-They suffer from tokenization overfragmentation [^2] See Appendix E.1 for a quantitative analysis.
-and thus require more inference steps due to the reliance on English-centric tokenizers and vocabulary, resulting in higher utility costs for non-English speakers Ahia et al.
+They suffer from tokenization overfragmentation (see Figure 2 for quantitative analysis) and thus require more inference steps due to the reliance on English-centric tokenizers and vocabulary, resulting in higher utility costs for non-English speakers Ahia et al.
 (2023); Petrov et al.
+(2023); Ali et al.
 (2024).
-Cross-lingual vocabulary adaptation (CVA) via vocabulary expansion has been proposed to adapt LLMs (including their tokenizers) to specific target languages.
+Cross-lingual vocabulary adaptation (CVA) via vocabulary expansion has been proposed to adapt LLMs (including their tokenizers) to specific target languages (Cui, Yang, and Yao, 2023; Fujii et al., 2024; Choi et al., 2024; Tejaswi, Gupta, and Choi, 2024; Mundra et al., 2024, inter alia).
 Vocabulary expansion approaches extend the vocabulary of a source model with tokens from a target language, followed by continual pre-training on target language data.
 A wide range of language-specific LLMs derived from an English-centric LLM such as Llama2 Touvron et al.
-(2023) have been made available following this approach, including Chinese Cui et al.
-(2023), Tamil Balachandran (2023), Portuguese Larcher et al.
+(2023) have been made available following this approach, including Chinese Cui, Yang, and Yao (2023), Tamil Balachandran (2023), Portuguese Larcher et al.
 (2023), and Japanese Fujii et al.
@@
 Vocabulary expansion improves inference speed but often assumes access to a substantial amount of target language data for adaptation.
-For example, Chinese and Tamil LLaMAs make use of 20 and 12GB of target language text, respectively Cui et al.
-(2023); Balachandran (2023).
-Given the guaranteed large number of model updates in high-resource settings, the embeddings of new tokens are often simply randomly initialized, followed by the standard continual pre-training with a causal language modeling (clm) objective Cui et al.
-(2023); Balachandran (2023); Larcher et al.
+For example, Chinese and Tamil Llamas make use of 20 and 12GB of target language text, respectively Cui, Yang, and Yao (2023); Balachandran (2023).
+Given the guaranteed large number of model updates in high-resource settings, the embeddings of new tokens are often randomly initialized, followed by the standard continual pre-training with a causal language modeling (clm) objective Cui, Yang, and Yao (2023); Balachandran (2023); Larcher et al.
 (2023); Choi et al.
@@
 However, it is not clear how effective this approach to vocabulary expansion is in low-resource settings.
-In this paper, we seek to (i) answer if this widely used adaptation approach under high-resource settings is as effective in low-resource settings; and (ii) identify the best possible vocabulary expansion strategies for language adaptation in low-resource settings, maintaining similar performance to the source models with faster inference (Figure 1).
+In this article, we seek to (i) answer if this widely used adaptation approach under high-resource settings is as effective in low-resource settings; and (ii) identify the best possible vocabulary expansion strategies for language adaptation in low-resource settings, while striving to maintain similar performance to the source model with faster inference (Figure 1).
 Our key contributions are as follows:
   • We present the first systematic study of vocabulary expansion-based adaptation of generative LLMs (i.e three English-centric models) in low-resource settings (i.e.
-assuming only 30K sentences, $\sim$0.01GB text data), on two generation tasks (i.e.
-machine translation and summarization) across ten typologically diverse languages.
+assuming only 30K sentences, $\sim$0.01GB text data or up to approximately 5M tokens), on two generation tasks (i.e.
+machine translation and summarization) and two classification tasks (i.e.
+multiple-choice reading comprehension and general knowledge and reasoning) across ten typologically diverse languages.
   • Our results show that the popular vocabulary expansion approach in high-resource settings (i.e.
 random initialization and fine-tuning the full model with a clm objective) is not always optimal in low-resource settings.
-  • We find that vocabulary initialization approaches that use heuristics information from the source and target tokinizers are more effective.
+  • We find that target parameter initialization approaches that use heuristics information from the source and target tokenizers are more effective.
 Furthermore, fine-tuning the top and bottom two layers of the LLM using a multi-token prediction objective Gloeckle et al.
@@
 Finally, using a short input sequence length by splitting longer text into multiple sentences allows for a larger number of model updates, mitigating underfitting.
-2 Related Work
-Text Overfragmentation.
+  • To better understand holistic aspects of vocabulary expansion in low-resource settings, we conduct a range of analyses, including the extent of source (English) knowledge retention and a direct comparison with vocabulary replacement.
+2 Background
+2.1 Text Overfragmentation
 LLMs tend to overfragment text in underrepresented languages Rust et al.
@@
 (2023); Petrov et al.
-(2023), slower inference Hofmann et al.
-(2022); Sun et al.
+(2023), slower inference Hofmann, Schuetze, and Pierrehumbert (2022); Sun et al.
 (2023); Petrov et al.
@@
 (2023); Fujii et al.
+(2023); Ali et al.
 (2024).
-Cross-lingual Vocabulary Adaptation.
-CVA methods have been proposed for adapting base LLMs to specific target languages for improving downstream performance and inference speed Cui et al.
-(2023); Balachandran (2023); Fujii et al.
-(2024); Yamaguchi et al.
+Figure 2: Average number of tokens on the FLORES-200 dev set across languages and models.
+To quantify this phenomenon, we calculate the average number of tokens on the FLORES-200 NLLB Team et al.
+(2022) dev set across our target languages and models (Figure 2).
+Following the established assumption that lower average token counts indicate more efficient tokenization Ahia et al.
+(2023), [^2] This assumption is valid specifically within the context of the FLORES-200 parallel machine translation corpus.
+our analysis reveals two key observations: (1) Burmese, Sinhala, and Telugu demonstrate the most severe overfragmentation among our ten target languages, requiring at least 6.8x more tokens than English across all models, and (2) English consistently exhibits the most efficient tokenization.
+2.2 Cross-lingual Vocabulary Adaptation
+CVA methods have been proposed for adapting base LLMs to specific target languages for improving downstream performance and inference speed Cui, Yang, and Yao (2023); Balachandran (2023); Fujii et al.
+(2024); Yamaguchi, Villavicencio, and Aletras (2024).
+CVA via vocabulary expansion incorporates new tokens into the source vocabulary Balachandran (2023); Larcher et al.
+(2023); Pipatanakul et al.
+(2023); Lin et al.
+(2024); Cui, Yang, and Yao (2023); Kim, Choi, and Jeong (2024); Fujii et al.
+(2024); Choi et al.
+(2024); Nguyen et al.
+(2024); Tejaswi, Gupta, and Choi (2024); Mundra et al.
 (2024).
-CVA via vocabulary expansion incorporates new tokens into the source vocabulary Chau et al.
-(2020); Cui et al.
-(2023); Balachandran (2023); Larcher et al.
-(2023); Fujii et al.
-(2024); Liu et al.
+CVA with vocabulary replacement replaces the entire or partial source vocabulary with a new one from the target Ostendorff and Rehm (2023); Csaki et al.
+(2023); Da Dalt et al.
+(2024); Remy et al.
+(2024); Yamaguchi, Villavicencio, and Aletras (2024); Dobler and de Melo (2024); Cahyawijaya et al.
 (2024).
-CVA with vocabulary replacement replaces the entire source vocabulary with a new one from the target Minixhofer et al.
-(2022); Dobler and de Melo (2023); Ostendorff and Rehm (2023); Downey et al.
+More recent methods include a hypernetwork for tokenizer transfer Minixhofer, Ponti, and Vulić (2024) and adapters for vocabulary alignment Han et al.
 (2025).
-More recent methods extend only the language modeling head Hong et al.
-(2024).
-CVA is followed by continual pre-training on target language data, often called language adaptive pre-training – LAPT Chau et al.
-(2020).
+CVA is typically followed by continual pre-training on target language data, often called language adaptive pre-training – LAPT Chau, Lin, and Smith (2020).
 Vocabulary expansion is widely used for developing language-specific generative LLMs from a source model, e.g.
-Chinese Cui et al.
-(2023) and Tamil Balachandran (2023) LLaMAs, and Swallow for Japanese Fujii et al.
+Chinese Cui, Yang, and Yao (2023) and Tamil Balachandran (2023) Llamas, and Swallow for Japanese Fujii et al.
 (2024).
@@
 This might not be feasible in low-resource settings with limited target language data or compute.
-To the best of our knowledge, our work is the first to investigate vocabulary expansion for efficient LLM inference in extremely low-resource settings by assuming access to a small amount of target language data.
+To the best of our knowledge, our work is the first to investigate vocabulary expansion for efficient decoder-based LLM inference in extremely low-resource settings by assuming access to a small amount of target language data (30K sentences).
 3 Problem Statement
-Let $\mathcal{M}_{\text{s}}$ be a source LLM with a vocabulary $\mathcal{V}_{\text{s}}$.
-Given an additional target vocabulary $\mathcal{V}_{\text{new}}$, the aim is to learn a model $\mathcal{M}_{\text{t}}$ with the identical architecture as $\mathcal{M}_{\text{s}}$ that supports a target vocabulary $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$.
-$\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{t}}$ have their corresponding tokenizer, $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
-First, a target language-specific auxiliary tokenizer $\mathcal{T}_{\text{aux}}$ is trained on target language data $\mathcal{D}$.
-$\mathcal{T}_{\text{t}}$ is then derived from $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{aux}}$ along with their corresponding vocabularies $\mathcal{V}_{\text{s}}$ and $\mathcal{V}_{\text{aux}}$ by incorporating top $k$ frequent tokens ($\mathcal{V}_{\text{new}}$) that are in $\mathcal{V}_{\text{aux}}\backslash(\mathcal{V}_{\text{s}}\cap\mathcal{V}_{\text{aux}})$ into $\mathcal{V}_{\text{s}}$, resulting in $\mathcal{V}_{\text{t}}$.
-Subsequently, $\mathcal{M}_{\text{t}}$ is initialized with the weights of $\mathcal{M}_{\text{s}}$.
-Its embedding and output layer matrices are expanded such that the former is of dimensionality $|\mathcal{V}_{\text{t}}|\times H_{\text{t}}$ and the latter $H_{\text{t}}\times|\mathcal{V}_{\text{t}}|$, where $H_{\text{t}}$ is the hidden dimensionality of $\mathcal{M}_{\text{t}}$.
-Each new token representation is initialized by applying a target vocabulary initialization method (§4).
-If the weights of both matrices are not tied, they are initialized separately.
-Finally, $\mathcal{M}_{\text{t}}$ is adapted to the target language by training it on $\mathcal{D}$ using a clm objective.
-4 Target Vocabulary Initialization
-After vocabulary expansion, the new embeddings should be initialized.
+The goal in this article is to expand the vocabulary of a source LLM to effectively support a target language in low-resource settings.
+This involves transitioning from a source model $\mathcal{M}_{\text{s}}$ (with vocabulary $\mathcal{V}_{\text{s}}$ and tokenizer $\mathcal{T}_{\text{s}}$) to a model $\mathcal{M}_{\text{t}}$ that supports an expanded target vocabulary $\mathcal{V}_{\text{t}}$ and tokenizer $\mathcal{T}_{\text{t}}$.
+More specifically, given $\mathcal{M}_{\text{s}}$, $\mathcal{V}_{\text{s}}$, $\mathcal{T}_{\text{s}}$, and target language data $\mathcal{D}$, $\mathcal{M}_{\text{t}}$ is constructed as follows:
+  1. 1.
+    Crafting the Auxiliary Tokenizer: A target language-specific auxiliary tokenizer is firstly trained on $\mathcal{D}$.
+This tokenizer is built from scratch using $\mathcal{D}$, allowing it to capture language-specific nuances.
+It comes with its own vocabulary, $\mathcal{V}_{\text{aux}}$.
+  2. 2.
+    Constructing the Target Vocabulary and Tokenizer:
+    1. (a)
+      The new tokens $\mathcal{V}_{\text{new}}$ are identified by taking the top $k$ most frequent tokens from $\mathcal{V}_{\text{aux}}$ that are not already present in $\mathcal{V}_{\text{s}}$ (i.e.
+$\mathcal{V}_{\text{new}}=\text{top }k\in\mathbb{N}\text{ tokens from }\mathcal{V}_{\text{aux}}\setminus(\mathcal{V}_{\text{s}}\cap\mathcal{V}_{\text{aux}})$).
+    2. (b)
+      The target vocabulary $\mathcal{V}_{\text{t}}$ is then formed by combining the original source vocabulary $\mathcal{V}_{\text{s}}$ with these newly identified tokens: $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$.
+    3. (c)
+      The target tokenizer $\mathcal{T}_{\text{t}}$ is subsequently derived to operate on this expanded $\mathcal{V}_{\text{t}}$.
+  3. 3.
+    Initializing and Adapting the Target Model:
+    1. (a)
+      The target model $\mathcal{M}_{\text{t}}$ begins as a copy of $\mathcal{M}_{\text{s}}$, inheriting its identical architecture and pre-trained weights.
+    2. (b)
+      Its embedding and output layer matrices are then expanded to accommodate the larger $\mathcal{V}_{\text{t}}$.
+Specifically, the dimensionality of the embedding layer becomes $|\mathcal{V}_{\text{t}}|\times H_{\text{t}}$ and that of the output layer becomes $H_{\text{t}}\times|\mathcal{V}_{\text{t}}|$, where $H_{\text{t}}$ is the hidden dimensionality of $\mathcal{M}_{\text{t}}$.
+    3. (c)
+      Each representation for a new token (i.e.
+tokens in $\mathcal{V}_{\text{new}}$) is initialized using a target parameter initialization method (detailed in §4).
+If the weights of both embeddings and language modeling head (i.e.
+output layer) are not tied, they are initialized separately.
+[^3] We follow the original configuration of the source model regarding weight tying for the embeddings and output layer to preserve its original behavior as closely as possible.
+    4. (d)
+      $\mathcal{M}_{\text{t}}$ undergoes continual pre-training (i.e. LAPT) on the target language data $\mathcal{D}$ using a causal language modeling (clm) objective.
+4 Target Parameter Initialization
+After vocabulary expansion (step 2 in §3), the new embeddings should be initialized.
 Our aim is to investigate the effectiveness and robustness of different initialization approaches in low-resource settings.
 For that purpose, we evaluate: (1) random initialization; (2) initialization based on auxiliary models; and (3) heuristic-based initialization.
+[^4] Our primary focus in the subsequent discussion is on embedding initialization for simplicity.
+However, if the weights of the embeddings and the language modeling head are not tied, the language modeling head should also be initialized independently, using the identical process applied to the embeddings.
 4.1 Random Initialization
-We randomly initialize new tokens by sampling from $\mathcal{N}(\mu,\sigma^{2})$, where $\mu$ and $\sigma$ are the mean and standard deviation of token embeddings of $\mathcal{M}_{\text{s}}$ (Random).
-This is the most simple and common approach when adapting English-centric LLMs to a target language via vocabulary expansion in high-resource settings Cui et al.
-(2023); Balachandran (2023); Larcher et al.
+For new tokens, we randomly initialize the weights of their embeddings by sampling from $\mathcal{N}(\mu,\sigma^{2})$.
+Here, $\mu$ and $\sigma$ are the mean and standard deviation of the token embeddings from $\mathcal{M}_{\text{s}}$ (Random).
+This is the most simple and common approach when adapting English-centric LLMs to a target language via vocabulary expansion in high-resource settings Cui, Yang, and Yao (2023); Balachandran (2023); Larcher et al.
 (2023); Choi et al.
@@
 (2017), for initialization.
-[^3] Due to resource constraints, we only report results using FOCUS as a representative initialization method that relies on auxiliary embeddings since it outperforms other similar methods such as WECHSEL Minixhofer et al.
-(2022) for vocabulary replacement.
+[^5] Due to resource constraints, we only report results using FOCUS as a representative initialization method that relies on auxiliary embeddings since it outperforms other similar methods such as WECHSEL Minixhofer, Paischer, and Rekabsaz (2022) for vocabulary replacement.
 The main assumption is that semantic transfer of embeddings should result in better initialization over Random.
 We apply FOCUS by tokenizing $\mathcal{D}$ using $\mathcal{T}_{\text{t}}$ and train a fastText model for each language.
-[^4] Preliminary experiments with FOCUS using off-the-shelf pre-trained word-level fastText models yielded lower performance (see Appendix E.2).
+[^6] Preliminary experiments with FOCUS using off-the-shelf pre-trained word-level fastText models yielded lower performance.
 4.3 Heuristic-based Initialization
@@
 Motivated by this, we evaluate a set of heuristic-based initialization methods that do not rely on any external data or model and can be applied to any language.
-Mean.
-A straightforward approach is to initialize each new token in $\mathcal{V}_{\text{new}}$ with the average of its corresponding source tokens obtained by using $\mathcal{T}_{\text{s}}$ Yao et al.
+Mean
+A straightforward approach is to initialize the weights of each new token in $\mathcal{V}_{\text{new}}$ by averaging the weights of their corresponding source tokens, which are identified using $\mathcal{T}_{\text{s}}$ Yao et al.
 (2021).
-Koto et al.
-(2021) and Gee et al.
+Koto, Lau, and Baldwin (2021) and Gee et al.
 (2022) have also followed a similar approach for vocabulary replacement in domain adaptation.
-Merge.
+More recently, Tejaswi, Gupta, and Choi (2024) and Mundra et al.
+(2024) have demonstrated the effectiveness of this mean initialization for vocabulary expansion under high-resource settings (i.e.
+at least 200M tokens and 2.5B tokens, respectively).
+However, our study specifically investigates low-resource settings with only 30K sentences, equivalent to at most 5M tokens.
+Merge
 Mean uses solely $\mathcal{T}_{\text{s}}$, which might produce subtokens from $\mathcal{V}_{\text{s}}$ that are not semantically related to the new target token.
 To overcome this issue, we propose using merge rules from $\mathcal{T}_{\text{t}}$ to effectively initialize $\mathcal{V}_{\text{new}}$.
-Merge rules describe how $\mathcal{T}_{\text{t}}$ can combine two subtokens into one, e.g.
-(‘a’, ‘t’) $\rightarrow$ ‘at’.
-Therefore, each new token in $\mathcal{V}_{\text{new}}$ can be decomposed into several existing subtokens from $\mathcal{V}_{\text{s}}$.
-We hypothesize that such subtokens can be more semantically related to the new token than those in Mean as they use information from $\mathcal{T}_{\text{t}}$.
-We first identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$.
-Subsequently, we generate a mapping from each new target token to source subtokens using the identified merge rules.
-Finally, for each new token, we compute the hierarchical mean of the associated source subtoken embeddings using the mapping information.
-Algorithm 1 in Appendix §A describes our method.
-Align.
-The tokenization of a given word can change dynamically depending on the context, e.g.
-‘the’ can be tokenized as (‘th’, ’e’) or (‘t’, ‘he’).
-Likewise, a new token in $\mathcal{V}_{\text{t}}$ is not always tokenized in the same way by $\mathcal{T}_{\text{s}}$ when operating on a sequence rather than a single token, which is the case for Mean.
+Merge rules describe how $\mathcal{T}_{\text{t}}$ can combine two subtokens into one.
+Through these rules, each new token in $\mathcal{V}_{\text{new}}$ can be decomposed into several existing subtokens from $\mathcal{V}_{\text{s}}$.
+For instance, consider a new token: ‘superherohype’ in $\mathcal{V}_{\text{new}}$.
+According to the merge rules in $\mathcal{T}_{\text{t}}$, it can first be decomposed into (‘superhero’, ‘hype’).
+If ‘superhero‘ is also a new token, it can be further decomposed into (‘super’, ‘hero’).
+This process continues until all constituent parts are tokens from $\mathcal{V}_{\text{s}}$ (in this case, ’super’, ’hero’, and ’hype’).
+We hypothesize that such hierarchically derived subtokens from $\mathcal{V}_{\text{s}}$ are more semantically related to the new token than those obtained by simple averaging, as they leverage the specific tokenization information embedded in $\mathcal{T}_{\text{t}}$.
+The initialization process is as follows:
+  1. 1.
+    We identify all merge rules in $\mathcal{T}_{\text{t}}$ that result in a token found in $\mathcal{V}_{\text{new}}$.
+  2. 2.
+    Using these identified merge rules, we generate a hierarchical mapping for each new target token down to its constituent source subtokens in $\mathcal{V}_{\text{s}}$.
+  3. 3.
+    For each new token, we compute the hierarchical mean of the embeddings of its associated source subtokens, guided by the mapping information.
+Align
+Initializing the weights of new tokens in $\mathcal{V}_{\text{t}}$ by simply averaging the weights of their constituent subtokens from $\mathcal{T}_{\text{s}}$ as in Mean can be suboptimal.
+This naive approach does not account for how tokenization might change in a full sequence, as opposed to a single token.
+Consider the word: ‘_cup’ and a new token: ’_cu’.777‘_’ stands for a whitespace.
+While Mean averages the weights of ‘_c’ and ‘u’ from $\mathcal{T}_{\text{s}}$ to initialize that of ’_cu’, this can be suboptimal.
+If $\mathcal{T}_{\text{s}}$ tokenizes ‘_cup’ as ‘_c’ + ‘up’, but $\mathcal{T}_{\text{t}}$ tokenizes it as ‘_cu’ + ‘p’, ‘_cu’ functions as a new, distinct subword unit.
+Its weight initialization should reflect this new role, perhaps focusing more on relevant overlapping components like ‘_c’.
 To obtain a more fine-grained semantic representation of a new token, we propose token alignment initialization that leverages mapping information between tokens tokenized with $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$ and the mapping frequency (i.e.
@@
 This process allows us to consider different tokenization variants for the initialization of a given token.
-For each sentence $x\in\mathcal{D}$, we first tokenize it using $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
-We then compare the two tokenized versions and generate a mapping list of tuples for each new token in $x$, e.g.
-a token $t_{5}$ in $\mathcal{V}_{\text{t}}$ is tokenized as [($t_{1}$, $t_{2}$), ($t_{3}$, $t_{4}$)] in $\mathcal{V}_{\text{s}}$.
-Subsequently, we concatenate all mapping lists of tuples in $\mathcal{D}$ for each new token $t\in\mathcal{V}_{\text{new}}$.
-For each $t$, we then construct a unique set of its constituent tuples, i.e.
-$\bm{T}_{t}=(\bm{t}_{1},\bm{t}_{2},\bm{t}_{3},\dots)$, and the corresponding frequency list $\bm{F}_{t}=(f_{\bm{t}_{1}},f_{\bm{t}_{2}},f_{\bm{t}_{3}},\dots)$, given the concatenated list in the previous step.
-Finally, we use $\bm{T}_{t}$ and $\bm{F}_{t}$ to aggregate the mapping information for token $t$ and generate its representation by computing $\sum_{\bm{t}\in\bm{T}_{t}}{\left[f_{\bm{t}}\frac{1}{|\bm{t}|}\sum_{t^{\prime}\in\bm{t}}{\bm{e}_{t^{\prime}}}\right]}$, where $\bm{e}_{t^{\prime}}$ is the vector representation (embedding or LM head) of the subtoken $t^{\prime}\in\mathcal{V}_{s}$.
+Specifically, the weights of each new token $t\in\mathcal{V}_{\text{new}}$ are initialized as follows.
+  1. 1.
+    Generate Mappings from $\mathcal{D}$:
+    1. (a)
+      For each sentence $x\in\mathcal{D}$, we first tokenize it using both $\mathcal{T}_{\text{s}}$ and $\mathcal{T}_{\text{t}}$.
+    2. (b)
+      We then compare these two tokenized versions to identify how each new token $t$ in $\mathcal{T}_{\text{t}}$’s output corresponds to sequences of subtokens from $\mathcal{T}_{\text{s}}$’s output.
+These correspondences are stored as a list of tuples, e.g.
+a token $t$ from $\mathcal{V}_{\text{t}}$ might map to $(t_{1},t_{2}),(t_{3},t_{4})]$ from $\mathcal{V}_{\text{s}}$ in different contexts.
+  2. 2.
+    Construct Unique Mappings and Frequencies:
+    1. (a)
+      We concatenate all such mapping lists from $\mathcal{D}$ for each new token $t$.
+    2. (b)
+      From this combined list, we construct a unique set of constituent tuples for $t$, denoted as $\mathbf{T}_{t}=(\mathbf{t}_{1},\mathbf{t}_{2},\mathbf{t}_{3},\dots)$, where each $\mathbf{t}_{i}$ is a sequence of subtokens from $\mathcal{V}_{\text{s}}$.
+    3. (c)
+      We also create a corresponding frequency list $\mathbf{F}_{t}=(f_{\mathbf{t}_{1}},f_{\mathbf{t}_{2}},f_{\mathbf{t}_{3}},\dots)$, indicating how often each unique tuple $\mathbf{t}_{i}$ appeared in the mappings for token $t$ on $\mathcal{D}$.
+  3. 3.
+    Compute Initial Representation: Finally, using $\mathbf{T}_{t}$ and $\mathbf{F}_{t}$, we aggregate the mapping information to generate the representation for token $t$ by computing $\sum_{\mathbf{t}\in\mathbf{T}_{t}}{\left[f_{\mathbf{t}}\frac{1}{|\mathbf{t}|}\sum_{t^{\prime}\in\mathbf{t}}{\mathbf{e}_{t^{\prime}}}\right]}$.
+Here, $\mathbf{e}_{t^{\prime}}$ is the embedding of the subtoken $t^{\prime}\in\mathcal{V}_{\text{s}}$, $f_{\mathbf{t}}$ is the frequency of mapping $\mathbf{t}$, and $|\mathbf{t}|$ is the number of subtokens in $\mathbf{t}$.
 5 Training Strategy
-Continual pre-training on $\mathcal{D}$ is an integral part of vocabulary expansion to improve the alignment of the newly initialized embeddings.
+Continual pre-training on $\mathcal{D}$ (i.e.
+LAPT) is an integral part of vocabulary expansion to improve the alignment of the newly initialized embeddings.
 To this end, we explore: (i) the training procedure, (ii) the objective function, and (iii) the input sequence length.
-Procedure.
-We use by default low-rank adaptation (LoRAHu et al.
-(2022) applied to all linear layers, following previous work on vocabulary expansion in high-resource settings Cui et al.
-(2023); Balachandran (2023); Abbasi et al.
+Training Procedure
+  • LoRA: We use by default low-rank adaptation (LoRAHu et al.
+(2022) applied to all linear layers, following previous work on vocabulary expansion in high-resource settings Cui, Yang, and Yao (2023); Balachandran (2023); Abbasi et al.
 (2023); Choi et al.
 (2024).
-We also consider the two-stage tuning process (2-stage) Cui et al.
-(2023), where only the embeddings and LM head are first updated, followed by tuning the LoRA modules.
+  • 2-stage: We also consider the two-stage tuning process (2-stage) Cui, Yang, and Yao (2023), where only the embeddings and language modeling head are first updated, followed by tuning the LoRA modules.
 This process can help minimize the risk of overfitting to the initial embedding state, which might be suboptimal Downey et al.
 (2023).
-Alternatively, we train only the top and bottom two layers (2x2 LS) following Remy et al.
+  • 2x2 LS: We train only the top and bottom two layers (2x2 LS) of the model, following Remy et al.
 (2024).
-This approach aims to calibrate only the parts closely related to the encoding and decoding of the target language, minimizing changes to the base model.
-Note that none of these approaches have ever been compared to the standard LoRA or against each other.
-We tune the embeddings and LM head in each case.
-Objective.
-We first evaluate clm, which has been used by previous work in high-resource settings (Cui et al., 2023; Balachandran, 2023; Larcher et al., 2023, inter alia), as a default training objective.
-We also consider a multi-token prediction (mtp) objective Gloeckle et al.
+This calibrates only the parts closely related to the encoding and decoding of the target language Wendler et al.
+(2024); Tang et al.
+(2024), minimizing changes to the source model.
+Note that none of these approaches have ever been compared to the standard LoRA or against each other. We tune the embeddings and language modeling head in each case.
+Objective Function
+  • clm: We first evaluate a causal language modeling objective (clm), which has been used by previous work in high-resource settings (Cui, Yang, and Yao, 2023; Balachandran, 2023; Larcher et al., 2023, inter alia), as a default training objective.
+  • mtp: We also consider a multi-token prediction (mtp) objective Gloeckle et al.
 (2024), where the model must predict multiple consecutive tokens at each timestep rather than a single token.
@@
 However, it has yet to be explored for continual pre-training for cross-lingual transfer.
-We experiment with one additional LM head (i.e.
-predicting two consecutive tokens at a time) and initialize the additional weights with those from the original LM head.
-[^5] We find in our preliminary analysis that random initialization does not work well (Appendix E.3).
-Sequence Length.
-Finally, we shorten the default input sequence length from 2,048 to 512, thereby increasing the number of training batches.
+We experiment with one additional language modeling head (i.e.
+predicting two consecutive tokens at a time) and initialize the additional weights with those from the original language modeling head.
+[^8] We find in our preliminary analysis that random initialization does not work well.
+Input Sequence Length
+We shorten the default input sequence length from 2,048 to 512, thereby increasing the number of training batches.
 We hypothesize that this is critical in low-resource settings, as a small number of model updates could be prone to underfitting.
@@
 6.1 Source Models
-We use Llama2 (7B) Touvron et al.
-(2023), an English-centric model, with its $\mathcal{T}_{\text{s}}$ based on byte-fallback Byte Pair Encoding (BPE) Sennrich et al.
-(2016) and $|\mathcal{V}_{\text{s}}|$ set to 32K in the experiments.
-We also use Llama3 (8B) Dubey et al.
-(2024) and Gemma2 (9B) Riviere et al.
+We use Llama2 7B Touvron et al.
+(2023), an English-centric non-instruction-tuned model, with its $\mathcal{T}_{\text{s}}$ based on byte-fallback Byte Pair Encoding (BPE) Sennrich, Haddow, and Birch (2016) and $|\mathcal{V}_{\text{s}}|$ set to 32K in the experiments.
+We also use Llama3 8B Dubey et al.
+(2024) and Gemma2 9B Riviere et al.
 (2024) for consistency in our analysis.
@@
 128K and 256K, respectively.
+Note that Llama2 and Llama3 have untied embedding and language modeling head weights, while the original configuration of Gemma2 includes weight tying.
 6.2 Target Languages and Data
@@
 This includes German (Indo-European) and Swahili (Niger–Congo) for the Latin script, and Arabic (Afroasiatic), Burmese (Sino-Tibetan), Greek (Indo-European), Hindi (Indo-European), Japanese (Japonic), Sinhala (Indo-European), Telugu (Dravidian), and Thai (Kra–Dai) for non-Latin scripts.
-We select these languages because of the availability of downstream task datasets with the same task formulation across languages.
+We select these languages because of the availability of downstream task datasets with the same task formulation across languages, with a particular focus on generation tasks.
 To simulate a realistic low-resource adaptation scenario, we follow Yong et al.
 (2023) in using 30K sentences per language.
+With $|\mathcal{D}|$ set to 30K, this equates to up to approximately 5M tokens.
 These sentences are randomly sampled from their language-specific subcorpus of CC-100 Conneau et al.
 (2020).
+Note that previous work on vocabulary expansion for generative LLMs Tejaswi, Gupta, and Choi (2024); Mundra et al.
+(2024) uses at least 200M tokens and 2.5B tokens respectively, which is at least 40 times larger than our training budget.
+[^9] For additional context, the BabyLM challenge Warstadt et al.
+(2023); Hu et al.
+(2024), where participants must pre-train a model from scratch under low-resource settings, utilizes training budgets of up to 100M words, motivated by the observation that children are exposed to fewer than 100M words by 13 years of age.
+While the BabyLM challenge is a challenging task, our setting, with access to less than 5M tokens, represents an even more extreme low-resource scenario, posing a distinct challenge for effective target language adaptation.
+6.3 Baselines
+We use the following two methods as our baselines:
+  1. 1.
+    Source: We use the off-the-shelf source base (i.e.
+non-instruction-tuned) model $\mathcal{M}_{\text{s}}$ without any adaptation, following Tejaswi, Gupta, and Choi (2024).
+This provides a crucial reference point, allowing us to quantify the inherent performance in the target language before language-specific tuning, and thereby clearly measure specific gains from vocabulary expansion and LAPT.
+  2. 2.
+    CPT-Only: We continue pre-training (CPT) Source on $\mathcal{D}$, retaining its original vocabulary $\mathcal{V}_{\text{s}}$.
+This differs from vocabulary expansion approaches, which expand the original vocabulary $\mathcal{V}_{\text{s}}$ to include new terms $\mathcal{V}_{\text{t}}=\mathcal{V}_{\text{s}}\cup\mathcal{V}_{\text{new}}$ as mentioned in §3.
+It is important to emphasize that these baselines do not offer any inference speedups.
+6.4 Evaluation
+Tasks
+We use both generation and classification target language tasks to evaluate each approach.
+For generation tasks, we use (1) English-to-target machine translation (mt) using FLORES-200 (NLLB Team et al., 2022) and (2) summarization (sum) including German MLSUM Scialom et al.
+(2020), GreekSUM Evdaimon et al.
+(2024), and XL-Sum Hasan et al.
+(2021) for the rest.
+Note that sum performance is not directly comparable between different languages as the data does not match across languages.
+For classification tasks, we employ multiple-choice reading comprehension (mc) using Belebele (Bandarkar et al., 2024) and Global MMLU (gmmlu(Singh et al., 2025) as a general knowledge and reasoning benchmark.
+Note that gmmlu does not support Burmese and Thai.
+Number of Samples
+Following Ahia et al.
+(2023), we use 500 random samples for generation tasks (sum and mt).
+mc and gmmlu use their full test sets for evaluation.
+Specifically, mc has 900 samples per language subset, while gmmlu has 14K samples.
+Prompt Templates
+For sum prompt templates, we translate the English templates from Ahia et al.
+(2023) using a machine translation API, as in Yong et al.
+(2023).
+For mt, we create an English template and then translate it into each target language.
+For mc and gmmlu, we follow the default template provided by Hugging Face LightEval Habib et al.
+(2023).
+The complete prompt templates are listed in Table 12 in the Appendix.
+6.5 Evaluation Metrics
+Task Performance
+We use accuracy for mc and gmmlu, chrF Popović (2015) for mt, and ROUGE-L Lin (2004) for sum. In the analysis (§7.3), we also use BLEURT Sellam, Das, and Parikh (2020) as an auxiliary metric for sum.
+We report average zero-shot performance across five different runs for the generation tasks, namely sum and mt.
+For the classification tasks, we report single-run three-shot performance for mc and five-shot performance for gmmlu as these tasks are deterministically evaluated with temperature set to zero.
+Perplexity
+We report perplexity on 100K language-specific held-out CC-100 sentences as an auxiliary metric for evaluating model performance.
+Inference Efficiency
+We measure inference efficiency as the number of tokens generated per second Hong, Lee, and Cho (2024).
+6.6 Implementation Details
+Hyperparameters
+We set $|\mathcal{V}_{\text{aux}}|$ to 50K across languages and the number of new target tokens $|\mathcal{V}_{\text{new}}|$ to 100 by default.
+We investigate the effect of varying $|\mathcal{V}_{\text{new}}|$ in §7.4.
+For LAPT, we train each model for two epochs with a batch size of 8, a maximum learning rate of 1e-4, and a sequence length of 2,048.
+We set a LoRA rank to 8 following previous work Cui, Yang, and Yao (2023); Abbasi et al.
+(2023); Lin et al.
+(2024).
+Table 13 in the Appendix details the hyperparameter configurations employed during both the training and inference phases.
+To make a fair comparison, we do not conduct any parameter tuning and use the same ones across all approaches.
+For sum, we truncate an article whenever it exceeds the maximum prompt length of 4,096 to avoid the CUDA out-of-memory error.
+Evaluation Metrics Computation
+To compute ROUGE-L, we split sentences with an mT5 Xue et al.
+(2021) tokenizer as preprocessing following Maynez, Agrawal, and Gehrmann (2023) and subsequently call rouge_scorer [^10] https://github.com/csebuetnlp/xl-sum/tree/master/multilingual_rouge_scoring to compute the metric.
+We use BLEURT-20 to compute BLEURT.
+[^11] https://github.com/google-research/bleurt
+Libraries and Hardware
+For Llama2 and Gemma2, we train tokenizers using SentencePiece Kudo and Richardson (2018) and convert them into the Hugging Face Tokenizers Moi and Patry (2023) format.
+For Llama3, we train tokenizers using Hugging Face Tokenizers.
+We implement our models using PyTorch Paszke et al.
+(2019), Hugging Face Transformers Wolf et al.
+(2020) and PEFT Mangrulkar et al.
+(2022).
+We preprocess datasets with Hugging Face Datasets Lhoest et al.
+(2021).
+For evaluation, we use Hugging Face LightEval Habib et al.
+(2023).
+We use either four NVIDIA V100 (32GB) or a single A100 (80GB) for LAPT.
+Evaluation utilizes a single NVIDIA V100 (32GB) or A100 (80GB).
+Each analysis utilizes a consistent hardware configuration to ensure accurate measurement of inference efficiency.
+7 Results and Analysis
+7.1 Target Parameter Initialization
+We analyze the effect of different target parameter initialization methods (§4) on task performance and inference efficiency using Llama2 as source.
+7.1.1 Task Performance and Perplexity
+Table 1 show the performance of all methods on generation tasks, while Table 2 shows the corresponding perplexities on the held-out language-specific dataset.
+Performance on Generation Tasks
+Models initialized with Mean and Align generally exhibit performance comparable to or better than Source.
+Specifically, they outperform (by >> 2 points) or match (within 2 points) Source in 16 and 15 out of 20 cases, respectively.
+Mean shows a positive gain over Source in 10 cases, while Align achieves this in 12 cases.
+These results aligns with their perplexity scores (Table 2), where Align generally yields the lowest perplexity across languages, closely followed by Mean.
+While Merge also matches or surpasses Source performance in 16 cases (with 10 positive gains), it often shows the largest perplexities among the three heuristic-based initialization methods.
+We speculate that Merge might be less informative than these two methods, as it does not rely on surface information for initialization, resulting in slightly larger perplexity.
+Table 1: Mean performance on generation tasks (mt and sum) over five runs in low-resource settings (30K sentences) using Llama2 as source.
+Green shades indicate positive performance change over Source per language and task, respectively.
 Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
@@
 Source .08.03 360.0 .03.00 240.1 .25.01 250.1 .08.03 250.2 .12.03 390.2 .03.03 220.1 .05.00 260.1 .01.00 270.4 .06.00 210.2 .06.02 220.2
-LAPT .18.00 360.2 .11.00 240.3 .20.00 210.1 .21.00 250.3 .21.00 380.2 .08.00 230.2 .07.00 274.1 .12.00 300.1 .09.00 290.1 .12.00 230.2
-+Speedup Random .07.00 320.3 .05.00 170.3 .28.00 230.3 .13.00 200.1 .09.00 350.2 .18.00 220.2 .07.00 290.1 .13.00 290.1 .06.00 260.2 .07.00 190.2
+CPT-only .18.00 360.2 .11.00 240.3 .20.00 210.1 .21.00 250.3 .21.00 380.2 .08.00 230.2 .07.00 230.1 .12.00 300.1 .09.00 290.1 .12.00 230.2
++Speedup
+Random .07.00 320.3 .05.00 170.3 .28.00 230.3 .13.00 200.1 .09.00 350.2 .18.00 220.2 .07.00 290.1 .13.00 290.1 .06.00 260.2 .07.00 190.2
 FOCUS .07.00 320.2 .03.00 160.3 .28.00 220.4 .17.00 180.1 .10.00 340.2 .19.00 220.2 .05.00 290.3 .18.00 290.1 .05.00 260.2 .06.00 170.1
@@
 Align .06.00 330.1 .04.00 150.7 .26.01 210.2 .16.00 170.2 .13.00 350.2 .20.00 230.1 .08.00 300.3 .17.00 290.0 .06.00 270.2 .07.00 220.2
-Table 1: Mean performance on mt and sum tasks over five runs in low-resource settings (30K sentences) using Llama2 as source.
-The baselines without vocabulary expansion are in gray.
-Darker blue and red shades indicate higher positive and negative relative performance change over Source per language and task, respectively.
-6.3 Baselines
-We use the following two methods as our baselines: (1) the source model $\mathcal{M}_{\text{s}}$ off-the-shelf without any adaptation (Source); (2) we continue pre-training Source on $\mathcal{D}$ by keeping $\mathcal{V}_{\text{s}}$ (LAPT).
-These do not offer any inference speedups.
-6.4 Tasks
-We evaluate all models using two tasks per target language in a zero-shot setting: (1) English-to-target machine translation (mt) using FLORES-200 NLLB Team et al.
-(2022) and (2) summarization (sum) including German MLSUM Scialom et al.
-(2020), GreekSUM Evdaimon et al.
-(2023), and XL-Sum Hasan et al.
-(2021) for the rest.666sum performance is not directly comparable between different languages as the data does not match across languages.
-Following Ahia et al.
-(2023), we use 500 random samples.
-The prompt templates can be found in Appendix B.
-6.5 Evaluation Metrics
-Task Performance.
-We use chrF Popović (2015) for mt and ROUGE-L Lin (2004) for sum. We report average performance over five different runs throughout the experiments and analysis.
-Perplexity.
-We report perplexity on the 100K language-specific held-out sentences sampled from the CC-100 subcorpus as an auxiliary metric for evaluating model performance.
-Inference Efficiency.
-We measure inference efficiency as the number of tokens generated per second Hong et al. (2024).
-6.6 Implementation Details
-We set $|\mathcal{V}_{\text{aux}}|$ to 50K across languages and the number of new target tokens $k$ to 100 by default.
-We investigate the effect of varying $k$ in §7.4.
-For LAPT, we train each model for two epochs with a batch size of 8, a maximum learning rate of 1e-4, and a sequence length of 2,048.
-We set a LoRA rank to 8 following previous work Cui et al.
-(2023); Abbasi et al.
-(2023); Lin et al.
-(2024).
-7 Results and Analysis
-7.1 Target Vocabulary Initialization
-We analyze the effect of different target vocabulary initialization methods (§4) on task performance and inference efficiency using Llama2 as source.
+Table 2: Perplexity on language-specific held-out dataset using Llama2 as source.
+Note that results are not comparable between models with gray and others due to their difference in vocabulary.
+Bold and underlined indicate the best and second-best perplexities among adapted models for each language.
 Model ar my de el hi ja si sw te th
 Source 8.3 5.0 35.7 4.8 6.4 20.4 3.5 47.2 2.4 9.4
-LAPT 4.2 2.7 10.9 2.9 3.2 5.4 2.3 12.7 1.8 4.3
+CPT-only 4.2 2.7 10.9 2.9 3.2 5.4 2.3 12.7 1.8 4.3
 Random 11.9 11.9 12.0 7.1 8.3 15.0 8.7 13.9 7.1 8.8
@@
 Align 9.3 11.2 11.7 5.9 6.3 15.1 8.0 13.5 6.7 7.6
-Table 2: Perplexity on language-specific held-out dataset using Llama2 as source.
-Note that results are not comparable between models with gray and others due to their difference in vocabulary.
-Bold and underlined indicate the best and second-best perplexities among adapted models for each language.
+The baseline Random and FOCUS models perform similarly (within 2 points) to or better than Source in 14 cases, with 10 and 7 cases showing positive gains, respectively.
+These results place them among the least effective approaches.
+FOCUS, in particular, shows the fewest cases with positive gains and the worst perplexity in six languages (i.e.
+Burmese, German, Hindi, Sinhala, Telugu, and Thai).
+These findings suggest that sophisticated initialization methods do not always guarantee superior performance, possibly due to underfitting of an auxiliary embedding model.
+Further, the popular Random approach is not always optimal in low-resource scenarios.
+Thus, we conclude that models initialized with Mean and Align are more likely to perform competitively with Source on generation tasks.
+Analyzing performance by language and task, Mean and Align models show positive gains over Source in six and eight out of ten languages for mt, respectively.
+In cases where gains are not observed, any performance degradation is negligible (within 2 points).
+Notably, they substantially outperform Source in Greek, Japanese, and Swahili mt, with improvements ranging from 8 to 17 points.
+However, for sum, these adapted models often do not yield a performance gain over Source in the majority of the languages, except for Japanese, Sinhala, Swahili, and Telugu.
+In particular, Burmese and Greek show a substantial performance drop of up to 9 and 8 points, respectively.
+We hypothesize that sum may necessitate more training data than mt because generating longer text (up to 128 tokens, using long context) requires models to have strong generative capabilities in the target language.
+We later address this challenge in §7.2 by showing that these performance gaps can be drastically narrowed using our alternative training strategies, further contributing to the competitiveness of vocabulary expansion approaches to Source.
+Turning to CPT-only (i.e.
+continual pre-training without vocabulary expansion), it demonstrates strong performance, matching or improving upon Source in 17 out of 20 cases, a higher count by at least one case than achieved by the models with the heuristic-based initialization.
+A direct comparison against the best-performing Align [^12] It is best-performing because it generally achieves the lowest perplexities across languages and is the most likely among different initialization methods to provide a positive gain over Source.
+further confirms this.
+CPT-only outperforms Align in 10 out of 20 cases, whereas Align only wins in four.
+These results demonstrate the overall superiority of CPT-only over vocabulary expansion approaches.
+This trend aligns with previous work on CVA Downey et al.
+(2023); Yamaguchi, Villavicencio, and Aletras (2024) suggesting that CPT-only can often perform better than vocabulary expansion approaches in low-resource settings, possibly due to its reliance on robust and well-aligned original embeddings.
+Table 3: Mean performance on classification tasks (mc and gmmlu) in low-resource settings (30K sentences) using Llama2 as source.
+Green shades indicate positive performance change over Source per language and task, respectively.
+Note that gmmlu does not support Burmese and Thai.
+Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
+mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu
+Source .29 .29 .26 - .43 .39 .27 .28 .25 .28 .40 .33 .24 .27 .31 .28 .28 .27 .29 -
+CPT-only .30 .29 .22 - .42 .38 .29 .29 .28 .28 .39 .34 .23 .27 .33 .27 .24 .27 .28 -
++Speedup
+Random .28 .25 .22 - .42 .38 .22 .27 .27 .26 .35 .30 .22 .27 .29 .25 .28 .26 .28 -
+FOCUS .29 .26 .28 - .41 .38 .24 .26 .28 .26 .36 .31 .29 .26 .28 .25 .24 .27 .28 -
+Mean .28 .25 .23 - .42 .38 .30 .26 .27 .26 .38 .29 .28 .26 .29 .25 .23 .27 .30 -
+Merge .29 .25 .25 - .43 .38 .27 .28 .26 .28 .37 .29 .27 .26 .28 .25 .28 .26 .29 -
+Align .28 .25 .26 - .40 .37 .31 .26 .27 .27 .37 .28 .27 .27 .29 .25 .22 .27 .31 -
+Performance on Classification Tasks
+Table 3 presents the performance of all methods on classification tasks.
+A distinct trend emerges compared to generation tasks: adapted models rarely yield a positive performance gain over Source.
+While all vocabulary expansion approaches exhibit competitive or better performance than Source in at least 12 cases, their positive gains are limited to a maximum of four cases (22%) (Mean and Align).
+This is far fewer than the maximum of 12 cases (60%) observed on the generation tasks.
+Notably, CPT-only performs on par with or outperforms Source in 16 out of 18 cases.
+Nonetheless, this contrasts with its strong performance on generation tasks: it shows positive gains in only 6 out of 18 cases (33%), compared to 65% on generation tasks.
+These results suggest that while continual pre-training (i.e.
+LAPT) can offer benefits, its impact on discriminative tasks in low-resource settings is limited and differs from its effect on generation tasks.
+We hypothesize these differences arise because our adaptation process (LAPT) primarily optimizes a causal language modeling objective on unlabeled target language data $\mathcal{D}$.
+While this objective is highly beneficial for generative tasks requiring fluency and extended text production, it may not directly translate to immediate gains on classification tasks.
+Classification tasks, in contrast, often heavily depend on the inherent semantic and factual knowledge of a model, frequently requiring only a single token generation for prediction.
+Therefore, while previous work on vocabulary expansion with LAPT Tejaswi, Gupta, and Choi (2024); Cui, Yang, and Yao (2023); Balachandran (2023); Choi et al.
+(2024); Yamaguchi et al.
+(2025) often reports performance improvements in both generation and classification tasks in resource-rich settings, our findings suggest that under extremely low-resource settings (i.e.
+30K sentences per language, approximately up to 5M tokens), LAPT does not inherently guarantee an improvement in these discriminative capabilities.
+Table 4: Inference efficiency for: (a) generation tasks (mt and sum) over five runs; and (b) classification tasks (mc and gmmlu).
+We measure inference efficiency by the number of tokens generated per second (T/S).
+The Source row shows both raw T/S and baseline 1.00x speedup per language and task.
+Bold and underlined indicate the best and second-best speedups across tasks and models with the same base model for each language.
+(a) Generation tasks
 Model Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
@@
 Source 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x
-LAPT 1.01x 1.13x 0.98x 1.01x 0.97x 1.05x 1.00x 0.97x 1.00x 1.01x 0.98x 1.03x 1.00x 1.00x 0.98x 1.21x 0.98x 0.99x 0.98x 1.01x
+CPT-only 1.01x 1.13x 0.98x 1.01x 0.97x 1.05x 1.00x 0.97x 1.00x 1.01x 0.98x 1.03x 1.00x 1.00x 0.98x 1.21x 0.98x 0.99x 0.98x 1.01x
 Random 1.17x 1.88x 1.18x 1.48x 0.95x 0.97x 1.32x 1.79x 1.19x 1.57x 1.04x 1.13x 1.56x 2.54x 0.92x 1.27x 1.93x 4.26x 1.06x 1.55x
@@
 Align 1.10x 1.98x 1.08x 1.90x 0.95x 1.04x 1.37x 1.67x 1.15x 1.77x 1.09x 1.13x 1.47x 2.58x 0.93x 1.26x 1.84x 4.42x 1.08x 1.63x
-Table 3: Inference efficiency measured by the number of tokens generated per second (T/S).
-The first row of Source corresponds to its respective T/S for each language and task.
-Bold and underlined indicate the best and second-best speedups across tasks and models with the same base model for each language.
-Task Performance and Perplexity.
-Table 1 shows the performance of all methods, while Table 2 shows the corresponding perplexities on the held-out language-specific dataset.
-We first find that models adapted with Mean and Align show comparable or better performance than Source in the majority of the cases.
-Specifically, they outperform or perform the same as Source in 13 and 14 out of 20 cases, respectively.
-This result aligns with the perplexity results in Table 2, where Align tends to exhibit the lowest perplexity across languages, followed by Mean.
-Further, some performance gaps in the rest of the cases are within only a few points ($\leq$ 3).
-In contrast, Merge shows worse performance than Align and Mean, outperforming or performing the same as Source in 11 cases.
-We speculate that Merge might be less informative than these two methods, as it does not rely on surface information for initialization, resulting in lower performance.
-Random performs the same as or better than Source in 12 cases, while FOCUS achieves this in 11 cases—the worst of the adapted models along with Merge.
-This is also evident in the perplexity results, where FOCUS shows the worst perplexity in six languages (i.e.
-my, de, hi, si, te, and th).
-These results suggest that sophisticated initialization approaches do not always work well, and the popular Random is not the best choice in low-resource settings.
-Thus, we conclude that models adapted with Mean and Align are more likely to perform competitively to Source.
-Next, we analyze the results by language and task.
-We see that the adapted models, especially Mean and Align, generally perform competitively with Source within a few points of degradation across all languages and substantially outperform it in Greek, Japanese and Swahili mt, with up to 17 points of improvement in Japanese.
-However, they often do not perform competitively in Burmese, German, Greek, and Hindi sum.
-We hypothesize that sum might require more training data than mt to perform well as generating long text (up to 128 generated tokens and using long context) requires models to have strong generative capabilities in a target language.
-We later address this challenge in §7.2 by showing that these performance gaps can be drastically narrowed using our alternative training strategies, further contributing to the competitiveness of vocabulary expansion approaches to Source.
-Inference Efficiency.
+(b) Classification tasks
+Model Arabic Burmese German Greek Hindi Japanese Sinhala Swahili Telugu Thai
+mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu
+19.4 29.9 11.1 - 36.6 34.3 14.0 23.4 16.9 24.1 33.4 34.0 10.9 15.2 35.4 33.1 11.1 12.8 16.7 -
+Source 1.00x 1.00x 1.00x - 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x 1.00x -
+CPT-only 1.17x 0.98x 1.24x - 1.04x 0.97x 1.22x 0.92x 1.17x 0.99x 1.03x 0.93x 1.27x 1.10x 1.04x 1.02x 1.02x 1.11x 1.20x -
+Random 1.71x 1.12x 2.12x - 1.03x 1.01x 1.68x 1.29x 1.71x 1.25x 1.02x 0.99x 2.10x 1.92x 1.05x 1.01x 1.84x 2.10x 1.42x -
+FOCUS 1.71x 1.12x 2.13x - 0.99x 1.01x 1.68x 1.29x 1.71x 1.23x 1.02x 0.99x 2.11x 1.92x 1.05x 1.00x 1.84x 2.11x 1.42x -
+Mean 1.71x 1.10x 2.13x - 1.02x 1.01x 1.68x 1.29x 1.71x 1.25x 1.03x 1.00x 2.10x 1.92x 1.05x 1.03x 1.84x 2.11x 1.42x -
+Merge 1.71x 1.12x 2.12x - 1.02x 0.99x 1.68x 1.29x 1.71x 1.25x 1.02x 0.99x 2.10x 1.93x 1.06x 1.01x 1.84x 2.10x 1.42x -
+Align 1.71x 1.11x 2.12x - 1.03x 1.02x 1.68x 1.29x 1.71x 1.25x 1.03x 0.98x 2.10x 1.92x 1.06x 1.03x 1.84x 2.10x 1.42x -
+7.1.2 Inference Efficiency
 Table 4 shows the inference efficiency of adapted models across tasks and languages.
@@
 For mt, we generally see smaller differences between models since its input is mostly in English, and thus, only the output contributes to inference speedups.
-Looking at the results by language and script, we observe that non-Latin script languages, except for Japanese, show substantial speedups of up to 4.58x.
-Japanese, on the other hand, exhibits moderate inference speedups of up to 1.13x.
-This is possibly due to the inclusion of Japanese (0.1%) and Chinese [^7] Japanese shares some scripts (i.e.
-kanji) with Chinese.
-data (0.13%) in the Llama2 training corpus Touvron et al.
-(2023).
-For a similar reason, German, which accounts for the largest size among non-English languages in the corpus (0.17%), shows no speedups across tasks.
-Although Swahili is not included in the training corpus, it does not seem to benefit well from vocabulary expansion with Llama2, possibly due to its Latin script (91% of the pre-training data).
-Recommendation.
-Using Mean or Align can provide better downstream performance and inference speedups.
+Similarly, for classification tasks (mc and gmmlu), the observed speedups are generally similar across models.
+This is primarily because these tasks involve processing a long input context, and their output typically requires only a single token, limiting the room for producing the difference between different approaches.
+Examining speedups by language and script, most non-Latin script languages (Arabic, Burmese, Hindi, Sinhala, Telugu, and Thai) exhibit substantial inference speedups, reaching up to 4.58x.
+In contrast, Japanese shows only moderate speedups, up to 1.13x.
+This is likely due to its prior representation in the Llama2 training corpus, which includes Japanese (0.1%) and Chinese (0.13%) data Touvron et al.
+(2023), with Chinese sharing common scripts (i.e.
+kanji).
+Similarly, German shows negligible speedups across tasks.
+This can be attributed to its largest non-English language representation (0.17%) in the Llama2 training corpus, implying its vocabulary is already well-covered.
+As a Latin script language, German is also inherently well-tokenized by the primarily Latin-script-based vocabulary of Llama2.
+This further suggests that expanding the vocabulary of a source model with a small number of target language tokens (i.e.
+100 in Table 4) does not substantially accelerate inference if the language is already well-represented in the vocabulary of the base model.
+Figure 2 further supports this trend, as Japanese and German both show far less text tokenization overfragmentation for Llama2 compared to Burmese, Greek, Hindi, Sinhala, Telugu, and Thai.
+Despite not being explicitly included in the Llama2 training corpus, Swahili shows limited speedups.
+This might be because its Latin script aligns well with the predominant Latin script pre-training data (91%), meaning its tokenization is already relatively efficient even without explicit target vocabulary expansion.
+Recommendation
+Using Mean or Align can provide better downstream performance on generation tasks and inference speedups across tasks.
 Vocabulary expansion contributes well to inference speedups when a target language is not included in pre-training data and is written in non-Latin scripts.
-[^8] If inference speedups do not matter, we recommend using LAPT in general.
+[^13] If inference speedups do not matter, we recommend using CPT-only in general.
 Since it does not involve parameter updates from scratch, it can lead to more stable task performance in low-resource settings.
@@
 We analyze the effectiveness of each training strategy introduced in §5.
-Due to limited compute, we only use the best-performing Align as the vocabulary initialization method and experiment with six languages with the largest speedups (i.e.
-ar, my, el, hi, si, and te) in sum.
+Given that LAPT does not typically improve classification performance in low-resource settings, we focus on generation tasks for brevity.
+[^14] The corresponding classification results are presented in Table 14 in the Appendix.
+While some fluctuations are observed, we confirm that no training strategy offers consistent improvements across tasks and languages.
+Due to limited compute, we only use the best-performing Align as the target parameter initialization method and experiment with six languages with the largest speedups (i.e.
+Arabic, Burmese, Greek, Hindi, Sinhala, and Telugu) in sum.
 Table 5 lists the performance of the adapted models.
+Table 5: Mean performance of Align models on generation tasks over five runs with Llama2 as source. Green indicates positive performance change over Source.
 Arabic Burmese Greek Hindi Sinhala Telugu
@@
 Source .08.03 360.0 .03.00 240.1 .08.03 250.2 .12.03 390.2 .05.00 260.1 .06.00 210.2
-LAPT .18.00 360.2 .11.00 240.2 .21.00 250.3 .21.00 380.2 .07.00 274.0 .09.00 290.1
+CPT-only .18.00 360.2 .11.00 240.2 .21.00 250.3 .21.00 380.2 .07.00 230.1 .09.00 290.1
 LoRA clm+2048 .06.00 330.1 .04.00 150.7 .16.00 170.2 .13.00 350.2 .08.00 300.3 .06.00 270.2
@@
 mtp+512 .17.00 330.2 .16.00 260.2 .24.00 230.2 .15.00 360.1 .13.00 330.3 .12.00 290.1
-Table 4: Mean performance of Align models (five runs) with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
-Looking at the results by training procedure, we observe that 2x2 LS drastically improves performance across tasks and languages, achieving up to 10 points improvement from LoRA (Burmese mt and sum with clm+2048).
+Looking at the results by training procedure, we observe that 2x2 LS improves performance across tasks and languages.
+Gains range from 1 point (Hindi sum) to 10 points (Burmese mt) compared to LoRA with clm+2048 (our default approach in §7.1).
+The sole exception is Arabic sum, where performance remains the same as the baseline.
+Notably, 2x2 LS also helps to partially mitigate performance drops observed in sum for languages like Burmese and Greek.
+We hypothesize that 2x2 LS can reduce the risk of underfitting by focusing on calibrating only the parts closely related to the encoding and decoding of the target language.
 This suggests that the frequently used full LoRA approach in high-resource settings is not the best training approach in low-resource settings.
-Although 2-stage does not consistently provide gains over LoRA, especially for mt, it does not show substantial performance degradation either.
-Next, we look at the effectiveness of mtp.
-Although it does not always improve performance over clm, it boosts Arabic mt performance across models, especially with 2x2 LS.
-Similar to 2-stage, its performance degradation, if it exists, is generally within a few points ($\leq 3$).
-Finally, usign a short sequence length (512) works well across models and languages, showing improvements of up to 9 points (LoRA+mtp in Burmese and 2-stage+clm in Greek mt), confirming our hypothesis (§5) that increasing the number of model updates can help avoid underfitting.
-Recommendation.
+Although 2-stage does not consistently provide gains over LoRA, especially for mt, it does not lead to substantial performance degradation either, with a maximum drop of 3 points for Greek mt, compared to the default LoRA with clm+2048.
+Next, we examine the effectiveness of mtp.
+Although it does not consistently improve performance over clm, it notably boosts Arabic mt performance across models, particularly when used with 2x2 LS.
+Similar to the 2-stage approach, any performance degradation observed with mtp is minor, staying within 2 points.
+The only exceptions across all cases are Burmese and Hindi mt with 2x2 LS, and Telugu sum with 2-stage, where we observe moderate drops of 3 to 4 points.
+Based on these results, while mtp does not offer universal gains, its consistent benefit for a specific language like Arabic mt and its generally stable performance (i.e.
+avoiding substantial drops) across other tasks make it a valuable strategy, particularly when used with 2x2 LS.
+[^15] It is worth noting that mtp typically offers an additional benefit of improving inference efficiency through self-speculative decoding Gloeckle et al.
+(2024), which can be a huge advantage in practical deployments, though this aspect is not the primary focus of our analysis and beyond the scope of this article.
+Finally, using a short sequence length (512) works well across models and languages, showing improvements of up to 9 points (LoRA+mtp in Burmese and 2-stage+clm in Greek mt), confirming our hypothesis (§5) that increasing the number of model updates can help avoid underfitting.
+Recommendation
 2x2 LS generally leads to performance improvements.
-The short training sequence length of 512 can also contribute to performance improvements in low-resource settings.
-Finally, mtp sometimes improves performance with no major performance drop in other cases.
+The short training sequence length of 512 can also aid low-resource settings.
+While mtp does not always offer improvements, its notable boost for a certain language and its overall stable performance, with no substantial degradation elsewhere, make it a viable and beneficial strategy for adaptation in low-resource environments.
 7.3 Experiments with Other Source LLMs
@@
 We apply the 2x2 LS+mtp+512 training strategy and initialize models with either Mean or Align along with Random as a baseline.
-Due to resource constraints, we only experiment with Burmese, Sinhala, and Telugu in the remainder of the paper, as they are the worst fragmented languages of our target languages (see Appendix E.1).
+Due to resource constraints, we only experiment with Burmese, Sinhala, and Telugu in the remainder of the article, as they are the worst fragmented languages of our target languages (Figure 2).
 Table 6 shows the performance of adapted models using Llama3 and Gemma2 as source.
+Table 6: Mean performance and inference speedup with Llama3 and Gemma2 as source. Green indicates positive performance change over Source. The speedup ratio corresponds to Align.
 Burmese Sinhala Telugu
-Llama3 mt sum mt sum mt sum
-Source .09.00 280.0 .19.00 300.0 .24.00 290.0
-LAPT .17.00 280.1 .21.00 310.1 .30.00 290.1
-+Speedup Random .19.00 230.2 .14.00 310.2 .13.00 290.1
-Mean .19.00 230.2 .12.00 320.1 .12.00 290.2
-Align .21.00 280.3 .22.00 330.1 .31.00 310.1
-Speedup 2.60x 3.52x 2.36x 3.19x 2.15x 3.55x
+Llama3 mt sum mc gmmlu mt sum mc gmmlu mt sum mc gmmlu
+Source .09.00 280.0 .41 - .19.00 300.0 .50 .37 .24.00 290.0 .51 .40
+CPT-only .17.00 280.1 .31 - .21.00 310.1 .47 .36 .30.00 290.1 .42 .38
++Speedup Random .19.00 230.2 .28 - .14.00 310.2 .23 .25 .13.00 290.1 .27 .26
+Mean .19.00 230.2 .25 - .12.00 320.1 .23 .23 .12.00 290.2 .26 .25
+Align .21.00 280.3 .28 - .22.00 330.1 .30 .26 .31.00 310.1 .38 .31
+Speedup 2.60x 3.52x 2.30x - 2.36x 3.19x 1.98x 1.87x 2.15x 3.55x 1.96x 1.85x
+Burmese Sinhala Telugu
+Gemma2 mt sum mc gmmlu mt sum mc gmmlu mt sum mc gmmlu
+Source .21.01 340.1 .67 - .19.00 340.2 .71 .50 .31.01 300.2 .74 .57
+CPT-only .28.00 280.3 .61 - .29.00 340.1 .69 .48 .43.00 290.1 .71 .56
++Speedup Random .24.00 260.3 .46 - .26.00 330.1 .63 .43 .35.00 280.2 .65 .47
+Mean .25.00 250.3 .49 - .27.00 320.2 .63 .42 .33.00 280.1 .60 .43
+Align .25.00 250.2 .32 - .28.00 320.1 .63 .42 .32.00 290.1 .59 .43
+Speedup 1.52x 1.57x 1.51x - 1.26x 1.38x 1.31x 1.20x 1.07x 1.10x 1.06x 1.13x
+Table 7: Perplexity on language-specific held-out dataset using Llama3 and Gemma2 as source. Results are not comparable between models with gray and others due to their difference in vocabulary.
+(a) Llama3
+Model my si te
+Source 3.7 3.5 3.0
+CPT-only 1.9 2.0 1.8
+Random 6.9 7.1 6.1
+Mean 6.7 6.8 5.8
+Align 6.0 5.5 4.6
+(b) Gemma2
+Model my si te
+Source 48.0 59.9 51.9
+CPT-only 4.6 4.3 5.2
+Random 9.0 6.9 6.7
+Mean 9.0 6.9 6.6
+Align 9.0 6.9 6.6
+Overall, models adapted with Align generally either perform competitively or outperform their respective Llama3 and Gemma2 base models (Source) on both generation tasks.
+The only exception is Burmese sum with Gemma2, which shows a substantial 9-point drop.
+In contrast, models adapted with Mean, along with Random, often underperform Align when using Llama3 as the base model.
+Specifically, they underperform Align in Burmese sum, Sinhala and Telugu mt tasks, and across almost all classification tasks.
+This underperformance is attributed to underfitting, as evidenced by their higher perplexities compared to Align (see Table 7).
+A consistent challenge across vocabulary expansion approaches, including Align, is substantial performance degradation on classification tasks.
+This degradation ranges from an 8-point drop (Sinhala mc and gmmlu with Gemma2) to a significant 35-point drop (Burmese mc with Gemma2) for the best-performing Align.
+This issue is not unique to these approaches, as CPT-only also frequently exhibits moderate to substantial degradation (3 to 10 points) on classification tasks across both models.
+This further highlights the inherent challenges of LAPT for classification tasks in low-resource settings, as discussed in §7.1.1.
+We later showcase in §8.2 that these substantial drops can be largely mitigated using a post-hoc, training-free method, achieving performance within 3 and 5 points of the Gemma2 base model for mc and gmmlu, respectively.
+Beyond downstream performance, a core advantage of vocabulary expansion lies in inference efficiency.
+Align models, for instance, consistently exhibit inference speedups, reaching up to 3.52x for Llama3 and 1.57x for Gemma2.161616For context, speedups in the range of 1.3x to 3x are commonly reported for speculative decoding techniques Agrawal, Jeon, and Lee (2024); Timor et al.
+(2025); Huang, Guo, and Wang (2025).
+Thus, the observed speedups generally represent a meaningful gain in inference efficiency for LLMs.
+These results confirm the versatility of our approaches in optimizing inference speed across different base model architectures.
+However, we note a relatively moderate inference speedup of up to 1.13x for Telugu with Gemma2, despite its higher text fragmentation rate than Sinhala (Figure 2).
+This can be attributed to lower target token ratios across tasks for both input and output contexts (Figure 4), suggesting an underutilization of newly added tokens and highlighting potential avenues for improvement in the selection of new tokens for specific languages.
+Table 8: Mean sum performance in ROUGE-L and BLEURT over five runs. Green indicates positive performance change over Source.
+Burmese Sinhala Telugu
+Llama2 ROUGE-L BLEURT ROUGE-L BLEURT ROUGE-L BLEURT
+Source 240.1 .03.00 260.1 .19.00 210.2 .13.00
+CPT-only 240.3 .04.00 23.0.1 .25.00 290.1 .22.00
++Speedup Random 240.1 .07.00 320.2 .26.00 290.1 .23.00
+Mean 250.1 .06.00 320.1 .25.00 290.1 .23.01
+Align 260.2 .08.01 330.3 .26.01 290.1 .24.00
+Llama3
+Source 28.00 .06.00 30.00 .24.00 29.00 .22.00
+CPT-only 280.1 .06.00 310.1 .25.00 300.1 .29.00
++Speedup Random 230.2 .05.00 310.2 .27.00 290.1 .24.01
+Mean 230.2 .05.00 320.1 .29.01 290.2 .26.00
+Align 280.3 .07.00 330.1 .31.00 310.1 .28.00
 Gemma2
-Source .21.01 340.1 .19.00 340.2 .31.01 300.2
-LAPT .28.00 280.3 .29.00 340.1 .43.00 290.1
-+Speedup Random .24.00 260.3 .26.00 330.1 .35.00 280.2
-Mean .25.00 250.3 .27.00 320.2 .33.00 280.1
-Align .25.00 250.2 .28.00 320.1 .32.00 290.1
-Speedup 2.24x 1.53x 2.26x 1.53x 1.03x 1.25x
-Table 5: Mean performance and inference speedup over five runs with Llama3 and Gemma2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
-Overall, models adapted with Align generally perform competitively with Gemma2 Source in Sinhala and Telugu sum tasks and often outperform Llama3 Source in both tasks and Gemma2 Source in mt.
-Further, Align models exhibit substantial inference speedups of up to 3.55x and 2.26x for Llama3 and Gemma2, respectively.
-These results confirm the versatility of our strategies regardless of the type of base model.
-Although models adapted with Mean rival Source, they, along with Random, underperform it in Burmese sum and Sinhala and Telugu mt tasks when using Llama3 as source.
-This can be due to underfitting, i.e.
-their perplexities are higher than Align (Appendix E.4).
-Examining the results by source model and task, we observe that the adapted models with Llama3 and Gemma2 as source show substantially better mt performance than those with Llama2 despite their similar model size.
+Source 340.1 .10.00 340.2 .32.00 300.2 .30.00
+CPT-only 280.3 .10.00 340.1 .32.01 290.1 .26.00
++Speedup Random 260.3 .10.00 330.1 .34.01 280.2 .26.00
+Mean 250.3 .09.00 320.2 .33.00 280.1 .27.00
+Align 250.2 .10.00 320.1 .33.01 290.1 .28.00
+Finally, examining the results by source model and generation task, we observe that adapted Llama3 and Gemma2 models show substantially better mt performance than those with Llama2 despite their similar model size.
 In particular, the Telugu-adapted models initialized with Align with Llama3 and Gemma2 as source obtain 19 and 20 points better performance than those Llama2 counterparts.
@@
 However, sum does not seem to follow the trend, i.e.
-the adapted models with Llama3 and Gemma2 do not outperform those with Llama2.
-In fact, their semantic-level performances measured by BLEURT Sellam et al.
-(2020) improve up to 7 points from those with Llama2 (Appendix E.4).
+the adapted models with Llama3 and Gemma2 do not outperform those with Llama2 when evaluated with ROUGE-L.
+In fact, their semantic-level performances measured by BLEURT Sellam, Das, and Parikh (2020) improve up to 7 points from those with Llama2 (see Table 8).
 Thus, adapted models actually enjoy their base model capabilities to improve performance at the semantic level rather than the surface level measured by ROUGE-L.
-Figure 2: Downstream performance and inference speedup in (a) mt and (b) sum across different $|\mathcal{V}_{\text{new}}|$. Red and gray dotted lines denote Source and LAPT.
 7.4 Target Vocabulary Size
-The extent to which a target language benefits from inference speedups varies across languages, models, and tasks, as observed from Tables 3 and 5.
+The extent to which a target language benefits from inference speedups varies across languages, models, and tasks, as observed from Tables 4 and 6.
 Although larger $|\mathcal{V}_{\text{new}}|$ can lead to faster inference, it comes with the risk of underfitting in low-resource settings.
@@
 17,953 for Chinese Llama.
-For comprehensiveness, we employ all three source models and follow the same setup as in §7.3.
-Figure 2 shows the performance changes and corresponding inference speedups with respect to different $|\mathcal{V}_{\text{new}}|$.
+For comprehensiveness, we employ all three source models and follow the same setup as in §7.3.181818Table 15 in the Appendix provides a qualitative example of how tokenization in Burmese changes with respect to $|\mathcal{V}_{\text{new}}|$ using Llama2 as source.
+Figure 3: Downstream performance and inference speedup in (a) mt and (b) sum across different $|\mathcal{V}_{\text{new}}|$. Red and gray dotted lines denote Source and CPT-only.
+Figure 4: Average target token ratio in input ($\times$) and output ($\bullet$) with respect to $|\mathcal{V}_{\text{new}}|$ across models, languages, and tasks.
+7.4.1 Generation Tasks
+Figure 3 shows the performance changes and corresponding inference speedups for different $|\mathcal{V}_{\text{new}}|$ on generation tasks.
 We first observe that the larger $|\mathcal{V}_{\text{new}}|$, the worse the task performance across different source models, languages, and tasks in general.
@@
 $|\mathcal{V}_{\text{new}}|=100$ in our setup) for tasks like sum to be competitive with Source in task performance.
-[…]
-Recommendation.
+Next, we observe that the larger $|\mathcal{V}_{\text{new}}|$, the faster the inference up to around $|\mathcal{V}_{\text{new}}|=\text{1K}$ in all the cases.
+After that point, the inference speedups tend to plateau or sometimes deteriorate.
+This can be partially due to underfitting, which hinders a model from effectively using $\mathcal{V}_{\text{new}}$.
+Indeed, the target token ratio in output at $|\mathcal{V}_{\text{new}}|=\text{5K}$ drops by 59.7% (Burmese), 9.5% (Sinhala), and 30.0% (Telugu) from its peak at $|\mathcal{V}_{\text{new}}|=500$ or 1K when using Llama2 on sum (Figure 4).
+Therefore, although larger $|\mathcal{V}_{\text{new}}|$ can shorten prompt length in a target language, it does not always guarantee inference speedups unless a model is well-trained to handle $\mathcal{V}_{\text{new}}$.
+Figure 5: Downstream performance and inference speedup in (a) mc and (b) gmmlu across different $|\mathcal{V}_{\text{new}}|$. Red and gray dotted lines denote Source and CPT-only.
+7.4.2 Classification Tasks
+Figure 5 shows the performance changes and corresponding inference speedups with respect to different $|\mathcal{V}_{\text{new}}|$ on classification tasks.
+Unlike generation tasks, larger $|\mathcal{V}_{\text{new}}|$ does not always result in worse task performance.
+This is especially evident in Llama2 and Llama3 across tasks.
+Nonetheless, Gemma2 adapted models with Random often exhibit performance degradation as $|\mathcal{V}_{\text{new}}|$ becomes larger across tasks and languages, suggesting the importance of target parameter initialization.
+We hypothesize that the observed difference between classification and generation tasks stems from the fundamental nature of the task, as discussed in §7.1.1 and §7.3.
+Classification often involves predicting only a single token, which requires less extensive and nuanced generation capabilities compared to generation tasks, where models must produce multi-token sequences.
+Therefore, a model might not need to be as well-trained to effectively utilize $\mathcal{V}_{\text{new}}$ for inference in classification, as it does not generate sequences incorporating those new tokens.
+While performance degradation is substantial in Llama3 and Gemma2 (as observed in §7.3), classification tasks generally appear more robust to changes in $|\mathcal{V}_{\text{new}}|$.
+This also explains why inference speedups tend not to plateau in most classification scenarios, as larger $|\mathcal{V}_{\text{new}}|$ can effectively shorten prompt length in a target language (Table 15 for an qualitative example), leading to continued speedup gains without the performance trade-offs seen on generative tasks.
+Recommendation
 Setting $|\mathcal{V}_{\text{new}}|$ between 500 and 1K for mt and 100 for sum can be a suitable starting point to maintain competitive performance to Source while benefiting greatly from inference speedups under low-resource settings.
-8 Conclusion
-We investigated cross-lingual vocabulary expansion in low-resource settings across vocabulary initialization approaches and training strategies.
-Our extensive experiments reveal that a widely used approach in high-resource settings is not always optimal in low-resource settings.
-In contrast, models adapted by our alternative strategies achieve faster inference while rivaling their base models in task performance.
-We supplement our analysis with specific recommendations for effective vocabulary expansion.
-In future work, we will compare vocabulary expansion and vocabulary replacement for efficient LM inference in low-resource settings.
-Limitations
-Tokenizer.
-Heuristic-based target vocabulary initialization with Merge assumes the use of a BPE-based tokenizer which is a common choice in recent LLMs, e.g.
+While classification tasks are more robust to changes in $|\mathcal{V}_{\text{new}}|$, it is often preferable to utilize either CPT-only or Source for better performance if: (i) inference speedups are not the primary priority, and (ii) a post-hoc, training-free performance degradation mitigation technique, as discussed later in §8.2, is not applied.
+8 Discussion
+8.1 Source Knowledge Retention
+Previous work Tejaswi, Gupta, and Choi (2024); Mundra et al.
+(2024) has reported that vocabulary expansion followed by LAPT can lead to catastrophic forgetting of the original capabilities of a source model.
+We measure the extent to which adapted models in low-resource settings (30K sentences) suffer from this phenomenon by evaluating them on English-centric reading comprehension and general knowledge and reasoning benchmarks.
+To this end, we employ the English subset of Belebele (mc) and mmlu Hendrycks et al.
+(2021).
+We use all three source models and follow the same setup as in §7.3 and §7.4.
+However, for brevity, we evaluate only the best-performing Align initialization method.
+Table 9 presents the corresponding results on mc and mmlu.
+Table 9: Source knowledge retention evaluation on English benchmarks. Each language name indicates that a model has undergone LAPT on its corresponding target language data (30K sentences).
+Burmese Sinhala Telugu
+Llama2 mc mmlu mc mmlu mc mmlu
+Source .52 .46 .52 .46 .52 .46
+CPT-only .53 .47 .52 .47 .52 .46
+Align .38 .37 .42 .39 .41 .37
+Burmese Sinhala Telugu
+Llama3 mc mmlu mc mmlu mc mmlu
+Source .88 .67 .88 .67 .88 .67
+CPT-only .86 .66 .86 .66 .87 .66
+Align .75 .46 .65 .39 .78 .50
+Burmese Sinhala Telugu
+Gemma2 mc mmlu mc mmlu mc mmlu
+Source .91 .72 .91 .72 .91 .72
+CPT-only .90 .72 .91 .72 .90 .72
+Align .88 .67 .89 .66 .84 .63
+Overall, we observe varying degrees of source knowledge retention depending on the adaptation method.
+CPT-only (i.e.
+continual per-training without vocabulary expansion) generally preserves original source model capabilities, showing negligible performance drops of up to 2 points across different base models.
+In contrast, adapted models initialized with Align exhibit moderate to substantial performance degradation on these English tasks, with average drops ranging from 5.3 points for Gemma2, to 10 points for Llama2, and a substantial 18.7 points for Llama3.
+These results corroborate the findings in Mundra et al.
+(2024), which noted that performance degradation in source language tasks tends to be more severe during early LAPT stages with vocabulary expansion.
+This suggests that while LAPT with Align is effective for target language adaptation, especially on generation tasks, it comes with a considerable trade-off in the retention of the original source capabilities.
+8.2 Can we recover performance degradation from vocabulary expansion in low-resource settings without training?
+As discussed in §7, vocabulary expansion in low-resource settings offers inference speedups and largely retains or even improves performance on generation tasks.
+However, it leads to performance degradation in (i) target language classification tasks (§7.3 and §7.4) and (ii) source language capabilities (§8.1).
+This section investigates whether such performance degradation can be recovered without additional training.
+Indeed, our work, ElChat Yamaguchi et al.
+(2025), provides a post-hoc, training-free method to recover the original capabilities of adapted models.
+ElChat: A Post-hoc Method to Mitigate Catastrophic Forgetting
+ElChat Yamaguchi et al.
+(2025) is a post-hoc, training-free method designed to restore the original capabilities of an LLM after it has undergone vocabulary expansion and continual pre-training on target language data.
+The method requires access to an instruction-tuned model that has been further supervised fine-tuned on labeled conversational data, enabling the source base model (i.e.
+$\mathcal{M}_{\text{s}}$) to follow instructions.
+This prerequisite is readily met in practice, as frontier models like Llama3 and Gemma 2 commonly offer both base and instruction-tuned variants.
+ElChat addresses catastrophic forgetting of original capabilities by employing a strategic combination of two core steps: model merging and copying special token weights.
+This allows for robust recovery of degraded performance without incurring further training costs.
+[^19] For more technical details, please refer to Yamaguchi et al.
+(2025).
+Table 10: Gemma2 performance and inference speedup using ElChat, a post-hoc training-free method to mitigate catastrophic forgetting. Green indicates positive performance change over Source.
+(a) Target language tasks
+Burmese Sinhala Telugu
+mt sum mc gmmlu mt sum mc gmmlu mt sum mc gmmlu
+Source .21.01 340.1 .67 - .19.00 340.2 .71 .50 .31.01 300.2 .74 .57
+CPT-only .28.00 280.3 .61 - .29.00 340.1 .69 .48 .43.00 290.1 .71 .56
++Speedup Align .25.00 250.2 .32 - .28.00 320.1 .63 .42 .32.00 290.1 .59 .43
++ ElChat .30.00 330.1 .64 - .31.00 350.1 .73 .46 .41.00 310.1 .74 .52
+Speedup 1.52x 1.57x 1.51x - 1.26x 1.38x 1.31x 1.20x 1.07x 1.10x 1.06x 1.13x
+(ElChat) 1.53x 1.73x 1.48x - 1.20x 1.39x 1.30x 1.18x 1.09x 1.14x 1.06x 1.02x
+(b) English tasks
+Burmese Sinhala Telugu
+Gemma2 mc mmlu mc mmlu mc mmlu
+Source .91 .72 .91 .72 .91 .72
+CPT-only .90 .72 .91 .72 .90 .72
+Align .88 .67 .89 .66 .84 .63
++ ElChat .93 .72 .93 .71 .93 .71
+Results
+Constrained by resources, we apply ElChat to adapted Gemma2 models initialized with Align to evaluate its impact on both target language and English tasks.
+[^20] Gemma2 is chosen for its notable performance degradation in target language tasks (not only in classification tasks but also in Burmese sum) and as a representative of more recent models than Llama2 and Llama3, enabling a practical and meaningful comparison within our budget.
+We observe from Table 10 (a) that ElChat consistently improves performance across target language tasks.
+Specifically, for classification tasks, the performance drop relative to Source is substantially reduced within 5 points (Telugu gmmlu), a notable improvement from up to 35 points (Burmese mc) without ElChat.
+Furthermore, ElChat not only recovers degraded performance but also maintains the superior generative capabilities of the adapted models, even showing improvements in some cases (e.g.
+Sinhala sum).
+Crucially, these performance recoveries are achieved while largely retaining the inference speedups gained from vocabulary expansion.
+The effect of ElChat is even more pronounced in English tasks (Table 10 (b)).
+The adapted models almost fully recover their original source capabilities, matching Source performance with a negligible drop of at most 1 point.
+This highlights the effectiveness of using a post-hoc, training-free method to mitigate performance degradation from vocabulary expansion without requiring any additional training.
+Recommendation
+Using a post-hoc, training-free method like ElChat enables an adapted model to restore the original capabilities of the corresponding source model without additional training.
+For target language tasks, this can lead to enhanced performance while preserving the inference speedups gained from vocabulary expansion.
+8.3 Comparison with Vocabulary Replacement
+Intuitively, vocabulary replacement, which typically replaces the entire source vocabulary with a new one from a target language, necessitates a greater number of training tokens than vocabulary expansion.
+This is attributed to its substantially larger number of new parameters requiring alignment.
+For instance, Dagan, Synnaeve, and Roziere (2024) found that over 50 billion tokens were necessary to swap a tokenizer at no performance cost in their domain adaptation experiments.
+To illustrate the challenges of vocabulary replacement for target language adaptation in low-resource settings compared to vocabulary expansion, we conduct a brief comparative study using Gemma2 as the source model.
+Experimental Setup
+We first train a target tokenizer with a vocabulary size of 32K using target language data $\mathcal{D}$ and the same training configurations as the base Gemma2 tokenizer.
+For target parameter initialization, we consider Random, FOCUS, and Mean.
+We then continually pre-train each initialized model using the same approach as in Sections 7.3 and 7.4 (i.e.
+2x2 LS+mtp+512).
+Table 11: Vocabulary replacement performance and inference speedup on target language tasks using Gemma2 as source.
+Green indicates positive performance change over Source.
+The speedup ratio corresponds to Mean.
+Burmese Sinhala Telugu
+mt sum mc gmmlu mt sum mc gmmlu mt sum mc gmmlu
+Source .21.01 340.1 .67 - .19.00 340.2 .71 .50 .31.01 300.2 .74 .57
+CPT-only .28.00 280.3 .61 - .29.00 340.1 .69 .48 .43.00 290.1 .71 .56
++Speedup Random .00.00 00.0 .22 - .00.00 00.0 .27 .25 .00.00 10.1 .24 .23
+FOCUS .00.00 40.1 .22 - .02.00 120.3 .27 .27 .00.00 30.0 .23 .27
+Mean .06.00 160.1 .26 - .01.00 230.3 .27 .26 .04.00 120.3 .34 .26
+Speedup 1.14x 2.27x 2.95x - 1.01x 1.36x 2.70x 2.44x 0.89x 0.57x 1.77x 1.91x
+Results
+Table 11 presents the performance and inference speedups of models adapted using vocabulary replacement.
+Overall, these models exhibit notably poor performance across all initialization approaches and tasks in low-resource settings.
+Specifically, Random and FOCUS largely fail to perform well on generation tasks, yielding near-zero scores in mt and low scores of up to 12 points in sum.
+While Mean shows slight improvement over Random and FOCUS, its performance remains substantially lower than that of Source, CPT-only, and crucially, the vocabulary expansion counterpart (Table 6).
+For instance, in Telugu mc, the best vocabulary replacement performance (Mean at 34 points) is still far below Source (74) and vocabulary expansion with Mean (60).
+The only advantage of vocabulary replacement lies in its superior inference speedups in some tasks (e.g.
+Burmese mc at 2.95x, Sinhala mc at 2.70x, and Telugu gmmlu at 1.91x), which stems from its entirely optimized vocabulary for the target language.
+These results clearly demonstrate the significant challenges of effectively applying vocabulary replacement for target language adaptation under low-resource settings, when compared to vocabulary expansion.
+8.4 Limitations
+Target Language Classification Performance
+While we demonstrate that ElChat, a post-hoc and training-free method, helps adapted models recover degraded performance in target language classification tasks, a moderate performance drop can still persist, as observed in Telugu gmmlu (Table 10).
+Furthermore, ElChat is not applicable in scenarios where only base models are available as source for adaptation (i.e.
+lacking an instruction-tuned variant).
+These limitations represent ongoing challenges that extend beyond the immediate scope of this article.
+Comparison with Fully Multilingual LLMs
+Due to resource constraints, this article does not include a direct comparison with fully multilingual models like MaLA-500 Lin et al.
+(2024) and EMMA-500 Ji et al.
+(2024, 2025).
+Such a comparison would help contextualize our results.
+Model Size
+Our experiments use models of up to 9B parameters due to resource constraints.
+Scaling these experiments to larger LLMs is a valuable direction for future work to confirm the generalizability of our findings.
+Tokenizer
+Heuristic-based target parameter initialization with Merge assumes the use of a BPE-based tokenizer which is a common choice in recent LLMs, e.g.
 Gemma2, Llama3, Llama2, Mistral Jiang et al.
 (2023), inter alia.
-Experimenting with other tokenizers, such as Unigram Kudo (2018), is beyond the scope of this paper.
-Languages.
-Our work covers ten linguistically diverse languages, i.e.
-Arabic, Burmese, German, Greek, Hindi, Japanese, Sinhala, Swahili, Telugu, and Thai, following previous work on CVA that has also tested a similar number of languages.
-For instance, Downey et al.
-(2023) tested ten languages, and Minixhofer et al.
-(2022) tested eight languages.
-Experimenting with more languages is an interesting avenue for future work but out of the scope of this paper given our limited computing capacity.
-Model Size.
-Our experiments consider LLMs of 7B, 8B, and 9B parameters.
-Investigating the performance of vocabulary expansion with larger models would be valuable in future studies.
-Limited access to computational resources did not allow us to experiment with larger LLMs.
-Tasks.
-Our experiments consider two generation tasks, i.e.
-machine translation and summarization, following Hong et al.
-(2024).
-Nonetheless, investigating different types of downstream tasks (e.g.
-reasoning) would be an interesting direction but out of the scope of this paper, given our limited computation and limited language resources available.
-Acknowledgments
-We would like to thank Miles Williams, Huiyin Xue, and Constantinos Karouzos for the comments.
-We acknowledge IT Services at the University of Sheffield for the provision of services for high-performance computing.
-Appendix
-Appendix A Pseudo-code for Merge Rule-based Initialization
-Algorithm 1 describes the pseudo-code of vocabulary initialization using merge rules.
-Data: $\mathcal{V}_{\text{s}}$, $\mathcal{V}_{\text{new}}$, Merge rules of $\mathcal{T}_{\text{t}}$: $Merges$, Embeddings of $\mathcal{M}_{\text{t}}$: $\bm{E}$
-Result: Updated $\bm{E}$
-1
-/* Identify any merge rule in $\mathcal{T}_{t}$ whose resulting token is in $\mathcal{V}_{\text{new}}$/
-2 $Merges_{\text{new}}\leftarrow\text{List}[~{}]$;
-3 for $m\in Merges$ do
-4       $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
-5       $t\leftarrow t_{1}+t_{2}$;
-6       if $t\in\mathcal{V}_{\text{new}}$ then
-7             $AddToMerges_{\text{new}}(m)$;
-8
-9       end if
-10
-11 end for
-12
-/* Generate a mapping from a new token to source subtokens using $Merges_{\text{new}}$/
-13 $Maps\leftarrow\text{Dict}[t\in\mathcal{V}_{\text{new}},\text{List}[~{}]]$;
-14 while $|Merges_{\text{new}}|\neq 0$ do
-15       $UnmapMerges\leftarrow\text{List}[~{}]$;
-16       for $m\in Merges_{\text{new}}$ do
-17             $t_{1},t_{2}\leftarrow m.split(\text{``}~{}\text{''})$ “ ” );
-18             $t\leftarrow t_{1}+t_{2}$;
-19             if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
-20                   $AddToMaps[t]([t_{1},t_{2}])$;
-21
-22            else if $t_{1}\in\mathcal{V}_{\text{new}}\cap t_{2}\in\mathcal{V}_{\text{s}}$ then
-23                   if $Maps[t_{1}]\neq\text{List}[~{}]$ then
-24                         $AddToMaps[t]([Maps[t_{1}],t_{2}])$
-25                  else
-26                         $AddToUnmapMerges(m)$
-27                   end if
-28
-29            else if $t_{1}\in\mathcal{V}_{\text{s}}\cap t_{2}\in\mathcal{V}_{\text{new}}$ then
-30                   if $Maps[t_{2}]\neq\text{List}[~{}]$ then
-31                         $AddToMaps[t]([t_{1},Maps[t_{2}]])$
-32                  else
-33                         $AddToUnmapMerges(m)$
-34                   end if
-35
-36             else
-37                   if $Maps[t_{1}]\neq\text{List}[~{}]\cap Maps[t_{2}]\neq\text{List}[~{}]$ then
-38                         $AddToMaps[t]([Maps[t_{1}],Maps[t_{2}]])$
-39                  else
-40                         $AddToUnmapMerges(m)$
-41                   end if
-42
-43             end if
-44
-45       end for
-46
-47      if $|UnmapMerges|=0$ then
-48             break;
-49
-50       end if
-51      $Merges_{\text{new}}\leftarrow UnmapMerges$;
-52
-53 end while
-54
-/* Obtain vector representation for each new token using $Maps$/
-55 for $t\in\mathcal{V}_{\text{new}}$ do
-56       $\bm{E}[t]\leftarrow ComputeHierarchicalMean(\bm{E}[Maps[t]])$;
-57 end for
-58
-Algorithm 1 Merge Rule-based Initialization
-Appendix B Prompt Templates
-We translate the English prompt templates provided by Ahia et al.
-(2023) for sum with a machine translation API, following Yong et al.
-(2023).
-For mt, we formulate a task-specific English prompt, followed by machine translation for each language.
-Table 8 shows the prompt templates used in our evaluation.
-Table 8: Prompt template for each task and language.
-Appendix C Implementation Details
-C.1 Language Adaptive Pre-training
-Table 6 lists the hyperparameters in LAPT.
-C.2 Evaluation
-sum.
-To compute ROUGE-L, we split sentences with an mT5 Xue et al.
-(2021) tokenizer as preprocessing following Maynez et al.
-(2023) and subsequently call rouge_scorer [^10] https://github.com/google-research/google-research/tree/master/rouge to compute the metric.
-To compute BLEURT, we use BLEURT-20, which is available from https://github.com/google-research/bleurt.
-Hyperparameters.
-Table 7 lists the parameters used during evaluation.
-To make a fair comparison, we do not conduct any generation parameter tuning and use the same ones across all approaches.
-For sum, we truncate an article whenever it exceeds the maximum prompt length of 4,096 to avoid the CUDA out-of-memory error.
-C.3 Libraries and Hardware
-For Llama2 and Gemma2, we train tokenizers using SentencePiece Kudo and Richardson (2018) and convert them into the Hugging Face Tokenizers Moi and Patry (2023) format.
-For Llama3, we train tokenizers using Hugging Face Tokenizers.
-We implement our models using PyTorch Paszke et al.
-(2019), Hugging Face Transformers Wolf et al.
-(2020) and PEFT Mangrulkar et al.
-(2022).
-We preprocess datasets with Hugging Face Datasets Lhoest et al.
-(2021).
-For evaluation, we use Hugging Face LightEval [^11] https://github.com/huggingface/lighteval.
-We use a single NVIDIA A100 (80GB) for all experiments.
-C.4 Code and Models
-Our code and models are available on https://github.com/gucci-j/lowres-cve.
-Appendix D Licenses
-This study uses various publicly available models and datasets with different licenses, as detailed below, all of which permit their use for academic research.
-D.1 Models
-Llama2 is licensed under the Llama 2 Community License Agreement.
-[^12] https://ai.meta.com/llama/license/ Likewise, Llama3 is licensed under the Llama 3 Community License Agreement.
-[^13] https://llama.meta.com/llama3/license/ Gemma2 is licensed under the Gemma Terms of Use Agreement.
-[^14] https://ai.google.dev/gemma/terms
-Figure 3: Average number of tokens on the FLORES-200 dev set across languages and models.
-D.2 Datasets
-FLORES-200 is distributed under CC BY-SA 4.0.
-XL-Sum is licensed under CC BY-NC-SA 4.0, while MLSUM is distributed under an MIT license.
-GreekSUM is not made public, but the official script for reproduction is available on https://github.com/iakovosevdaimon/GreekSUM.
-Appendix E Results and Analysis
-E.1 Overfragmentation
-Figure 3 shows the average number of tokens on the FLORES-200 dev set across languages and models.
-Following the assumption by Ahia et al.
-(2023) that “language A is more efficiently tokenized than language B if it uses fewer tokens per sentence on average,” [^15] Note that this assumption is only applicable to the FLORES-200 dataset as it is a parallel machine translation corpus.
-we can clearly see that (1) Burmese, Sinhala, and Telugu are the worst overfragmented languages among the ten target languages across all models, and (2) English is the most efficiently tokenized language.
-E.2 Target Vocabulary Initialization
-Table 9 compares models adapted with FOCUS using Llama2 as source in Burmese, Sinhala, and Telugu.
-Subword-fastText is the exact same model as FOCUS in Table 1 and makes use of only $\mathcal{D}$ to train the fastText model.
-Word-fastText uses a pre-trained language-specific word-level fastText model provided by Grave et al.
-(2018), i.e.
-with access to an external model.
-We observe that models using a word-level pre-trained fastText model consistently underperform those using a subword-level fastText model except for sum in Sinhala.
-Therefore, we conclude that using a pre-trained word-level fastText model does not help improve performance for models adapted with FOCUS in low-resource settings.
-Burmese Sinhala Telugu
-Model mt sum mt sum mt sum
-Source .03.00 240.1 .05.00 260.1 .06.00 210.2
-LAPT .11.00 240.3 .07.00 274.1 .09.00 290.1
-FOCUS
-Subword-fastText .03.00 160.3 .05.00 290.3 .05.00 260.2
-Word-fastText .01.00 50.2 .04.00 290.3 .03.00 240.2
-Table 9: Mean performance of models initialized with different variants of FOCUS over five runs with Llama2 as Source.
-Darker blue and red indicate higher positive and negative relative performance change over Source.
-Arabic Greek Hindi
-Model mt sum mt sum mt sum
-Source .08.03 360.0 .08.03 250.2 .12.03 390.2
-LAPT .18.00 360.2 .21.00 250.3 .21.00 380.2
-LoRA mtp + 2048 .11.00 320.2 .14.00 200.2 .12.00 340.2
-mtp $\ominus$ Copy + 2048 .10.00 320.1 .15.00 200.1 .11.00 350.2
-2-stage mtp + 2048 .13.00 330.1 .19.00 210.1 .11.00 360.1
-mtp $\ominus$ Copy + 2048 .10.01 330.1 .17.00 220.2 .10.00 360.1
-2x2 LS mtp + 2048 .15.01 330.1 .23.00 230.2 .12.00 360.0
-mtp $\ominus$ Copy + 2048 .09.00 330.1 .15.00 210.1 .09.00 350.1
-Table 10: Mean performance of models initialized with Align over five runs with Llama2 as Source. Darker blue and red indicate higher positive and negative relative performance change over Source.
-Model my si te
-Source 3.7 3.5 3.0
-LAPT 1.9 2.0 1.8
-Random 6.9 7.1 6.1
-Mean 6.7 6.8 5.8
-Align 6.0 5.5 4.6
-(a) Llama3
-Model my si te
-Source 48.0 59.9 51.9
-LAPT 4.6 4.3 5.2
-Random 9.0 6.9 6.7
-Mean 9.0 6.9 6.6
-Align 9.0 6.9 6.6
-(b) Gemma2
-Table 11: Perplexity on language-specific held-out dataset using Llama3 and Gemma2 as source.
-Note that results are not comparable between models with gray and others due to their difference in vocabulary.
-E.3 Training Strategies
-Table 10 briefly compares mtp approaches with and without additional LM head initialization based on the original LM head weights in Arabic, Greek, and Hindi.
-$\ominus$ indicates that an additional LM head is randomly initialized.
-We can see that although models without additional LM head initialization based on their original LM head weights show similar performance to those counterparts in sum, they generally do not outperform those in mt.
-Therefore, we conclude that it is essential to initialize the weights of an additional LM head with those from the original LM head for effective continual pre-training under low-resource settings.
-Burmese Sinhala Telugu
-Llama2 ROUGE-L BLEURT ROUGE-L BLEURT ROUGE-L BLEURT
-Source 240.1 .03.00 260.1 .19.00 210.2 .13.00
-LAPT 240.3 .04.00 27.4.1 .25.00 290.1 .22.00
-+Speedup Random 240.1 .07.00 320.2 .26.00 290.1 .23.00
-Mean 250.1 .06.00 320.1 .25.00 290.1 .23.01
-Align 260.2 .08.01 330.3 .26.01 290.1 .24.00
-Llama3
-Source 28.00 .06.00 30.00 .24.00 29.00 .22.00
-LAPT 280.1 .06.00 310.1 .25.00 300.1 .29.00
-+Speedup Random 230.2 .05.00 310.2 .27.00 290.1 .24.01
-Mean 230.2 .05.00 320.1 .29.01 290.2 .26.00
-Align 280.3 .07.00 330.1 .31.00 310.1 .28.00
-Gemma2
-Source 340.1 .10.00 340.2 .32.00 300.2 .30.00
-LAPT 280.3 .10.00 340.1 .32.01 290.1 .26.00
-+Speedup Random 260.3 .10.00 330.1 .34.01 280.2 .26.00
-Mean 250.3 .09.00 320.2 .33.00 280.1 .27.00
-Align 250.2 .10.00 320.1 .33.01 290.1 .28.00
-Table 12: Mean sum performance in ROUGE-L and BLEURT over five runs. Darker blue and red indicate higher positive and negative relative performance change over Source.
-E.4 Experiments with Other Source LLMs
-Perplexity.
-Table 11 shows the perplexities measured on the language-specific held-out dataset using Llama3 and Gemma2 as source.
-ROUGE-L vs. BLEURT.
-Table 12 shows the sum performances measured by ROUGE-L and BLEURT across models when using the 2x2 LS + mtp + 512 strategy and the Align initialization.
-As discussed in §7.3, while performances at the surface level measured by ROUGE-L do not substantially change across different base models, those at the semantic level measured by BLEURT change, especially between Llama2 and Gemma2.
-E.5 Target Vocabulary Size
-Figure 4 shows the average target token ratio across models, tasks, and languages with respect to different $|\mathcal{V}_{\text{new}}|$.
-Llama2.
+Experimenting with other tokenizers, such as Unigram Kudo (2018), falls outside the scope of this article.
+9 Conclusion
+We investigated cross-lingual vocabulary expansion in low-resource settings across target parameter initialization approaches and training strategies.
+Our extensive experiments reveal that a widely used approach in high-resource settings is not always optimal in low-resource settings.
+In contrast, models adapted by our alternative strategies achieve faster inference while rivaling their base models in task performance, especially on generation tasks.
+We supplement our analysis with specific recommendations for effective vocabulary expansion.
 […]
-Llama3.
-We observe that the ratio of target tokens for outputs plateaus at around $|\mathcal{V}_{\text{new}}|=500$ in both tasks.
-After that, there are some drops in the target token ratio in mt.
-Interestingly, the ratio of target tokens for outputs is quite similar to that for inputs in sum.
-suggesting that adapted models effectively utilize newly added tokens in outputs.
-To further optimize inference efficiency, we hypothesize that adding target tokens with a larger number of characters might work.
-We leave this verification for future work.
-Gemma2.
+Details on Experimental Setup
+Table 12 lists prompt templates for all tasks, while Table 13 lists the hyperparameters used for both (a) LAPT and (b) inference.
 […]
-Moreover, the ratio of target tokens is far lower than that of Llama2 and Llama3, indicating that newly added tokens are underutilized in Gemma2.
-This could possibly be due to byte-fallback BPE used by Gemma2.
-Considering Llama3 (i) employs byte-level BPE, (ii) shows substantial speedups with vocabulary expansion in Table 5 and Figure 2, (iii) has also a large vocabulary size of 128K, and (iv) has similar fragmentation rates to Gemma2 for the three languages in Figure 3, vocabulary expansion might work better with byte-level BPE rather than byte-fallback BPE.
-We hypothesize that since byte-level BPE operates on a byte sequence rather than a raw text, it is, therefore, not affected by language scripts.
-This nature can help to integrate both a new vocabulary and merge rules with the corresponding source ones.
-We leave a thorough investigation of this hypothesis for future work.
-Figure 4: Average target token ratio in input ($\times$) and output ($\bullet$) with respect to $|\mathcal{V}_{\text{new}}|$ across models, languages, and tasks.
+Table 13: Hyperparameter configurations for LAPT and inference.
+(a) LAPT
+Hyperparameters Values
+Batch size 8
+Maximum number of training epochs 2
+Adam $\epsilon$ 1e-8
+Adam $\beta_{1}$ 0.9
+Adam $\beta_{2}$ 0.999
+Sequence length 2,048
+Learning rate 1e-4
+Learning rate scheduler cosine
+Warmup steps 100
+Weight decay 0.01
+Attention dropout 0.0
+Dropout 0.05
+LoRA rank $r$ 8
+LoRA dropout 0.05
+LoRA $\alpha$ 32
+(b) Inference
+Parameters Values
+Maximum prompt length 4,096
+Temperature 0.8
+Repetition penalty 1.1
+Top $k$ 40
+Top pp 0.9
+Beam width 5
+Sampling True
+Early stopping True
+Maximum number of generated tokens 128
+[…]
+Supplementary Results
+Table 14 lists the performance of the adapted models with different training strategies on classification tasks.
+Table 15 visualizes how tokenization changes with vocabulary expansion in Burmese using Llama2 as source.
+Table 14: Mean performance of Align models on classification tasks with Llama2 as source. Green indicates positive performance change over Source.
+Arabic Burmese Greek Hindi Sinhala Telugu
+Model mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu mc gmmlu
+Source .29 .29 .26 - .27 .28 .25 .28 .24 .27 .28 .27
+CPT-only .30 .29 .22 - .29 .29 .28 .28 .23 .27 .24 .27
+LoRA clm+2048 .28 .25 .26 - .31 .26 .27 .27 .27 .27 .22 .27
+mtp+2048 .29 .26 .28 - .26 .28 .28 .25 .28 .26 .27 .27
+clm+512 .28 .25 .24 - .33 .26 .28 .27 .27 .26 .27 .26
+mtp+512 .29 .28 .24 - .27 .27 .30 .26 .24 .25 .27 .27
+2-stage clm+2048 .30 .27 .24 - .26 .25 .27 .26 .27 .27 .27 .26
+mtp+2048 .29 .27 .22 - .23 .24 .28 .26 .29 .26 .27 .26
+clm+512 .25 .28 .22 - .29 .27 .27 .27 .27 .25 .26 .27
+mtp+512 .23 .28 .22 - .27 .28 .29 .26 .28 .25 .22 .27
+2x2 LS clm+2048 .29 .27 .22 - .29 .24 .28 .26 .27 .26 .28 .26
+mtp+2048 .30 .27 .26 - .32 .25 .28 .27 .28 .24 .29 .27
+clm+512 .31 .29 .22 - .28 .24 .28 .28 .27 .26 .28 .26
+mtp+512 .28 .27 .24 - .23 .23 .28 .26 .27 .26 .27 .26
+Table 15: Example of tokenization changes with vocabulary expansion in Burmese using Llama2 as source. ‘_’ stands for a whitespace.
+Acknowledgements.
+We thank the anonymous reviewers and Action Editor, Minlie Huang, for their constructive and detailed feedback.
+We are also grateful to Miles Williams, Huiyin Xue, and Constantinos Karouzos for their valuable feedback on the initial draft.
+We acknowledge (1) IT Services at the University of Sheffield for providing high-performance computing services, (2) the EuroHPC Joint Undertaking for awarding us access to MeluXina at LuxProvide, Luxembourg, and (3) the use of time on Tier 2 HPC facility JADE2, funded by the Engineering and Physical Sciences Research Council (EPSRC) (EP/T022205/1).
+AY is supported by EPSRC [grant number EP/W524360/1] and the Japan Student Services Organization (JASSO) Student Exchange Support Program (Graduate Scholarship for Degree Seeking Students).
+AV research is partly supported by MRC-FAPESP [AIM-Health] and CNPq [406926/2025-5].
+[…]